### v1

# GlossLM: Multilingual Pretraining for Low-Resource Interlinear Glossing

Abstract
A key aspect of language documentation is the creation of annotated text in a format such as interlinear glossed text (IGT), which captures fine-grained morphosyntactic analyses in a morpheme-by-morpheme format. Prior work has explored methods to automatically generate IGT in order to reduce the time cost of language analysis. However, many languages (particularly those requiring preservation) lack sufficient IGT data to train effective models, and crosslingual transfer has been proposed as a method to overcome this limitation.
We compile the largest existing corpus of IGT data from a variety of sources, covering over 450k examples across 1.8k languages, to enable research on crosslingual transfer and IGT generation. Then, we pretrain a large multilingual model on a portion of this corpus, and further finetune it to specific languages. Our model is competitive with state-of-the-art methods for segmented data and large monolingual datasets. Meanwhile, our model outperforms SOTA models on unsegmented text and small corpora by up to 6.6% morpheme accuracy, demonstrating the effectiveness of crosslingual transfer for low-resource languages.111 Our pretrained model and dataset are available at https://huggingface.co/lecslab/glosslm and https://huggingface.co/datasets/lecslab/glosslm-corpus-split, and will be made accessible through a web interface for use in language documentation efforts.
[…]
GlossLM: Multilingual Pretraining for Low-Resource Interlinear Glossing
[…]
1 Introduction
footnotetext: * Equal contribution
With nearly half of the world’s 7,000 languages considered endangered, communities of minoritized language speakers are working to preserve and revitalize their languages (Seifart et al., 2018). These efforts often involve the collection, analysis, and annotation of linguistic data. Annotated text is especially useful in the creation of reference texts (like dictionaries and grammars) and other archival materials, as well as in the development of language technologies including searchable digital text Blokland et al. (2019); Rijhwani et al. (2023) and computer-assisted educational tools Uibo et al. (2017); Chaudhary et al. (2023)
Figure 1: Components of interlinear gloss with an Arapaho sentence and English translation Cowell (2020). Blue boxes show transcriptions that are unsegmented (top) or segmented (bottom). Segmented text is split into morphemes which are aligned with the gloss labels shown in the green box. The task of automatic glossing uses some or all of the information in the gray box (transcription & translation) to generate the gloss line.
One prevalent form of linguistic annotation in language documentation projects is interlinear glossed text (IGT). IGT is a multi-line data format which includes (1) a transcription of speech in the language, (2) an aligned morpheme-by-morpheme description, and oftentimes (3) a free translation. An example of each of these components is illustrated in Figure 1. IGT is useful for illustrating morphosyntactic features of languages that other linguists or researchers may not be familiar with, and can also be used as examples for grammatical sketches or language learning textbooks, among other applications.
Creating IGT corpora is time-consuming and difficult, requiring documenters to perform linguistic analysis and extensive documentation simultaneously, often for languages with limited prior documentation. Research has found that computational tools can help accelerate the annotation process and overcome this bottleneck Palmer et al. (2009). In particular, models that take as input transcriptions and translations (which don’t require linguistic training to produce) and output glosses can rapidly bootstrap creation of complete IGT datasets.
The majority of prior work on automatic glossing focuses on training monolingual models that can predict IGT for a single language (Moeller and Hulden, 2018; McMillan-Major, 2020; Zhao et al., 2020). This approach has two major drawbacks. First, many language community members or linguists do not have the expertise, time, or resources to create special-purpose models, creating a high barrier of entry for adoption. Second, in the rare cases where expertise is available, there is often insufficient amounts of data to train effective monolingual models, with many IGT corpora having only a few thousand examples or less.
In this work, we aim to overcome the monolingual data bottleneck by creating a multilingual pretrained gloss model that can be further adapted to specific languages and gloss formats. Inspired by recent advances in massively multilingual language modeling, we compile various existing IGT corpora into a dataset containing over 450k IGT examples in 1.8k languages, making it the largest existing collection of digitized IGT. We then train our model, GlossLM, by continually pretraining ByT5 Xue et al. (2022), and evaluate it on seven typologically diverse languages. We find that our model is highly competitive with existing models, outperforming them broadly on the more challenging task of glossing unsegmented text (where words are not split into individual morphemes), as well as in certain cases where segmented text is provided. Furthermore, we observe that our model retains high performance on seen languages even without finetuning, indicating that it can be used as-is without specialization to particular languages.
2 Interlinear Glossed Text (IGT)
2.1 Description of IGT Format
IGT is a structured data format which presents a transcription of text in a language being studied, along with its gloss—aligned labels that mark each morpheme’s meaning and/or grammatical function—in a different language. Oftentimes, a free translation in a widely-spoken language is included. An IGT example for Arapaho is given in subsection 2.1 Cowell (2020), with glosses and translations in English:
[…]
nuhu’ tih-’eeneti-3i’ heneenei3oobei-3i’
this when.PAST-speak-3PL IC.tell.the.truth-3PL
[…]
This example is segmented, with morphemes separated by dashes. Each morpheme in the Arapaho sentence (e.g. tih) is directly aligned with glosses (separated by periods) that describe that morpheme’s function and/or meaning. Labels in all caps (e.g. PAST) are grammatical glosses; lowercase labels (e.g. speak) are lexical glosses.
IGT examples may also use unsegmented transcriptions, as in the Uspanteko example in subsection 2.1 Pixabaj et al. (2007). In this case, words and their labels are aligned, but no explicit alignment between morpheme labels and the original morphemes themselves is present, and thus the segmentation of words into morphemes is unclear:
[…]
o sey xtok rixoqiil
o sea COM-buscar E3S-esposa
[…]
IGT is highly valuable to documentation projects, and has also been used in the NLP context for the creation of morphological paradigms (Moeller et al., 2020), machine translation (Zhou et al., 2019), generating precision grammars (Bender et al., 2013), and other tools including POS taggers and dependency parsers (Georgi, 2016).
2.2 Challenges with Automatic Glossing
Effective glossing requires knowledge of the target language as well as linguistic understanding of morphological patterns. Furthermore, certain factors exist that make this task particularly difficult for automated systems.
Lack of segmented transcriptions: As in subsection 2.1, often only an unsegmented transcription is available. In these cases, an automated system cannot predict glosses morpheme-by-morpheme, but must instead learn to perform simultaneous segmentation and glossing.
Variance in glossing styles: While there are widely-used standards for glossing conventions, such as the Leipzig glossing rules Comrie et al. (2008), exact glossing styles may vary widely from documenter to documenter Chelliah et al. (2021). Documentation projects may have differences in label spelling (e.g. SING/SG/S to denote singular), formatting (e.g. underscores/periods for multi-word lexical glosses), and language specific labels (e.g. a linguist may want to specify different levels of an intensifier marker), among others.
Extremely low-data settings: When IGT is used for language documentation, the target languages of interest are likely those with very small amounts of digital data to begin with.
Given these challenges, we want our model to be easily adaptable to different languages with varying amounts of training data, as well as able to handle both segmented and unsegmented inputs.
3 Dataset
We compile and clean the largest IGT dataset from a variety of sources and languages. In total, our dataset contains over 450k IGT instances (from 250k unique sentences) in 1.8k languages, collected from six different IGT corpora. All sources are publicly available under the CC BY 4.0 License, allowing free use and redistribution, and we have confirmed with the creators of each source that our usage is within the intended use. We make our artifacts available under the same license.
3.1 Data Sources
Corpus Languages IGT instances
ODIN 936 83,661
SIGMORPHON 7 68,604
IMTVault 1,116 79,522
APICS 76 15,805
UraTyp 35 1,719
Guarani Corpus 1 803
Table 1: Number of unique examples and languages in each source corpus for the GlossLM dataset.
ODIN
The Online Dictionary of Interlinear Text (ODIN, Lewis and Xia 2010) is a large dataset of 158k IGT examples representing 1496 languages, compiled by scraping IGT from linguistics documents on the internet. We use the preprocessed version of ODIN by He et al. (2023), which discards languages with less than five IGT samples, resulting in 84k unique glossed sentences across 936 languages.
SIGMORPHON Shared Task
We use the IGT data from the 2023 SIGMORPHON Shared Task on Interlinear Glossing (Ginn et al., 2023). The data covers seven languages with diverse features and includes 69k glossed sentences. We use the shared task corpora as our primary evaluation sets, with the same splits as the shared task.
IMTVault
IMTVault (Nordhoff and Krämer, 2022) is a recent aggregation of IGT data extracted from code in books published by the Language Science Press. We use the 1.1 release (Nordhoff and Forkel, 2023) which includes 1116 languages and 80k examples.
APiCS
The Atlas of Pidgin and Creole Language Structures (APiCS) is a set of books detailing grammatical features of 76 pidgin and creole languages (Michaelis et al., 2013a, b). APiCS online provides interactive versions of the books, including 16k IGT examples.
UraTyp
UraTyp (Norvik et al., 2022) provides grammatical and typological information, collected from linguistic questionnaires on various languages. This includes a small number of IGT examples (1.7k) spanning 35 languages.
Guarani Corpus
The Guarani Corpus (Maria Luisa Zubizarreta, 2023) consists of 803 examples of IGT, representing fifteen stories, for Guarani, a Tupian language spoken in South America. We use Beautiful Soup [^2] https://pypi.org/project/beautifulsoup4/ to parse examples from HTML.
3.2 Preprocessing
In total we have 250k unique IGT instances in 1.8k languages. We use Hugging Face Datasets to store our corpus in a widely-used, publicly accessible format. Each row retains a unique ID and reference to its source, so examples can be easily cross-referenced.
If datasets explicitly indicate whether an IGT line is segmented, we use this value. Otherwise, we determine segmentation by checking if a line has any morpheme boundary markers (the hyphen “-”). For segmented words, we remove the segmentation markers to create an additional unsegmented version of the same example, for a total of 451k examples (206k segmented).
We run langid (Lui and Baldwin, 2012) on translations to verify the translation language labels, and leave the translation field blank if the predicted language did not match the language indicated by the original source. Finally, we pad any non-lexical punctuation with spaces and normalize spacing, as our experiments indicate that our models are sensitive to this formatting.
3.3 Language Coverage
Within our dataset, around 90% of examples have an associated Glottocode (Hammarström et al., 2023), amounting to 1,785 unique Glottocodes and over 150 language families represented. While it would be ideal to have a relatively balanced set across languages and language families, many languages only have a few lines of IGT available, and thus our dataset has a long tail distribution across languages. The language with the greatest representation by far is Arapaho (from the SIGMORPHON Shared Task dataset) with almost 98k IGT instances, making up about 20% of the entire dataset. Overall, 25% of languages have fewer than 5 IGT instances, 50% have fewer than 10, and 75% have fewer than 54. We include histograms for the distributions across languages and language families in Appendix A, as well as preliminary analysis of typological coverage using the Grambank database Skirgård et al. (2023) in Appendix B.
3.4 Target Languages
We select data from the SIGMORPHON Shared Task to use for evaluation and testing, as this data is highly consistent and has been evaluated on prior models. We use the same splits as the shared task (Table 2) and designate three languagues–Arapaho, Tsez, and Uspanteko–as in-domain languages, which are included in the pretraining corpus. We use Gitksan, Lezgi, Natugu, and Nyangbo as out-of-domain languages, which do not occur in the pretraining corpus, but are used to finetune models. All languages except Nyangbo include translations.
Finetuning
Language Pre-train Train Eval Test
Other languages 313790 - - -
In-domain languages
Arapaho (arp) 78,264 78,264 9,784 9,784
Tsez (ddo) 7116 7116 890 890
Uspanteko (usp) 19,584 19,584 464 1,266
Out-of-domain languages
Gitksan (git) - 163 84 74
Lezgi (lez) - 1,411 176 174
Natugu (ntu) - 1,582 198 198
Nyangbo (nyb) - 4,200 526 526
Table 2: Number of total (segmented and unsegmented) pretrain (for in-domain languages), train, evaluation, and test examples for the target languages.
4 GlossLM Model
Using our IGT dataset, we aimed to create a single multilingual pretrained model that could be easily adapted to documentation projects, both for known languages and unseen ones.
4.1 Architecture
We base GlossLM off of ByT5, a multilingual pretrained model using the T5 architecture Raffel et al. (2020). ByT5 operates on byte-level inputs, as opposed to word or subword tokens, making it easily adaptable to different scripts and unseen languages. We use the ByT5-base model (582M parameters), pretrained on the mC4 dataset Xue et al. (2021). We did not experiment with pretraining a randomly initialized model, as pretraining runs are expensive and we predict that the pretrained non-IGT base model serves as a better initialization.
4.2 IGT Pretraining
We continually pretrain ByT5 on two versions of the train split of our dataset. The first (GlossLMall) uses all the training data in our dataset. Our second model (GlossLMunseg) excludes any segmented data in any of our test languages, simulating scenarios where a linguist does not have access to a corpus of segmented glossed data in their target language. Then, for each version of GlossLM, we finetune on the target language of interest, again using all data in the target language for GlossLMall and only unsegmented data in the target language for GlossLMunseg.
As ByT5 is based on the T5 model, we cast the glossing task into a text-to-text format. We train the model using examples formatted with the following prompt:
Provide the glosses for the following transcription in <lang>.
Transcription in <lang>: <transcription>
Transcription segmented: <yes/no/unknown>
Translation in <metalang>: <translation>
Glosses:
Models are trained to output the gloss line following the above prompt input. Including translations has been shown to help in predicting glosses (Zhao et al., 2020; Ginn et al., 2023). For some data, a translation was not available ($\approx 3\%$ of the training data), in which case the translation line is omitted. We pretrain models using the hyperparameters given in Appendix C. We did not conduct hyperparameter search, only tuning the batch size, to fit in our GPUs, and epochs and early stopping, to ensure convergence.
5 Experiments
We evaluate the effectiveness of our pretrained model on automatic generation of IGT across the seven evaluation languages described in §3.4.
5.1 Experimental Settings
We compare models under two settings.
Segmented
In the segmented setting, words in the input transcriptions are divided into morphemes with segmentation boundaries. Models must learn to classify each morpheme with an appropriate gloss, making this an easier setting (with SOTA models achieving 80-90% accuracy).
Unsegmented
In the unsegmented setting, words in the input transcriptions are not segmented (but glosses are), requiring the model to jointly learn segmentation and glossing. Generally, this setting is far more difficult, with SOTA models achieving as low as 11% accuracy on small datasets.
5.2 GlossLM Experiments
We evaluate the GlossLM model in two variations. First, we directly evaluate the pretrained model on the in-domain languages without any additional finetuning, in order to analyze whether the model can retain high performance across individual languages despite the highly multilingual pretraining corpus. We refer to these models as GlossLMpre.
We then finetune the GlossLM pretrained model on the training dataset for each language individually, and label these runs as GlossLMft. We do so for both the in-domain languages, to focus the model on the target language, as well as the out-of-domain languages, allowing us to study the model’s adaptation to unseen languages. Finetuning used the same parameters as pretraining, but with 100 training epochs and early stopping (patience 3, start epoch 15) [^3] For Gitksan, due to the size of the training set, we set max epochs to 300 and patience to 15., and took anywhere from 20 minutes to around 4 hours for each language. Inference uses beam search with $n=3$ beams.
5.3 Baseline Experiments
We compare the GlossLM models with six baseline systems which include the state-of-the-art from prior work.
Top-choice
selects the most frequent label associated with each morpheme/word in the training data, and assigns “???” to unseen morphemes.
Token-class
treats the glossing task as a token classification problem, where the output vocabulary consists of the IGT morpheme or word-level labels. Each target language’s data is used to train a language-specifc Token-class model, which uses the RoBERTa architecture with default hyperparameters without any additional pretraining Liu et al. (2019). This was used as the baseline model for the SIGMORPHON 2023 Shared Task on interlinear glossing Ginn (2023).
Tü-CL Girrbach (2023b)
uses straight-through gradient estimation (Bengio et al., 2013) to induce latent, discrete segmentations of input texts, and predicted glosses using a multilayer perceptron.
CRF Okabe and Yvon (2023b)
is a hybrid CRF-neural system that is composed of an unsupervised neural alignment system (SimAlign, Jalili Sabet et al. 2020) to make use of lexemes in translations and a CRF sequence labeling system (Lost, Lavergne et al. 2011).
SigMoreFun (SMF) He et al. (2023)
also continually pretrain ByT5, though on a smaller amount of data (using the ODIN corpus). Data from all target languages were combined and shuffled for finetuning.
ByT5all and ByT5unseg
are ByT5 models Xue et al. (2022) finetuned on the target language data without any multilingual continual pretraining on IGT. We otherwise replicate the finetuning procedure described in subsection 5.2.
5.4 Evaluation
For evaluating predictions, we strip punctuation (except for within glosses). [^4] Because of this post-processing, our results for baselines are slightly different than what original sources report. We evaluate morpheme accuracy, which counts the number of correct morpheme glosses in the correct position. Hence, if a gloss is inserted or deleted, the following glosses will be incorrect. We also evaluate word accuracy, which counts the number of entire correct word glosses.
However, accuracy may sometimes be too strict of a measurement—especially for generative models—as minor character insertions/deletions in the label are penalized heavily. Thus, we also evaluate chrF++ (Popović, 2015), a character-level metric often used in machine translation. chrF++ measures the F1 score over character n-grams between the reference and predictions, and is robust to insertions and deletions, unlike accuracy.
Model Morpheme accuracy / Word accuracy (Segmented)
In-domain Out-of-domain
arp ddo usp Avg git lez ntu nyb Avg
Top-choice 83.2 / 74.0 78.5 / 64.4 79.7 / 72.9 80.5 / 70.4 51.1 / 29.7 62.2 / 54.4 78.4 / 68.1 72.5 / 63.8 66.1 / 54.0
Token-class 90.4 / 84.2 86.4 / 76.5 82.5 / 76.5 86.4 / 79.1 25.3 / 16.4 50.2 / 38.8 62.0 / 54.3 88.8 / 84.4 56.6 / 48.5
Tü-CL 90.7 / 84.6 91.2 / 85.1 84.5 / 78.5 88.8 / 82.7 50.2 / 26.6 84.9 / 77.6 91.7 / 86.0 91.4 / 87.9 79.6 / 69.5
CRF 90.4 / 84.2 91.9 / 85.6 84.4 / 78.9 88.9 / 82.9 52.4 / 33.6 84.7 / 77.5 91.1 / 86.6 88.8 / 84.4 79.3 / 70.5
SMF 80.1 / 79.4 78.2 / 82.8 73.2 / 75.7 77.2 / 79.3 12.7 / 20.6 47.8 / 56.4 64.0 / 75.7 85.4 / 82.7 52.5 / 58.9
ByT5all 88.7 / 83.2 93.5 / 89.9 86.3 / 82.7 89.5 / 85.3 2.2 / 3.6 72.5 / 69.7 83.4 / 82.2 90.7 / 89.2 62.2 / 61.2
GlossLMall, pre 89.3 / 84.2 91.7 / 88.3 84.1 / 81.0 88.4 / 84.5 3.6 / 9.1 3.6 / 1.8 4.9 / 9.8 2.9 / 3.0 3.8 / 5.9
GlossLMall, ft 90.1 / 85.0 92.8 / 89.3 86.4 / 84.5 89.8 / 86.3 28.9 / 34.9 74.7 / 71.3 86.0 / 81.5 90.7 / 87.7 70.1 / 68.9
Model Morpheme accuracy / Word accuracy (Unsegmented)
In-domain Out-of-domain
arp ddo usp Avg git lez ntu nyb Avg
Top-choice 27.9 / 56.9 15.2 / 64.1 43.6 / 60.4 28.9 / 60.5 3.6 / 16.9 20.1 / 58.2 12.7 / 55.1 72.3 / 76.7 27.2 / 51.7
Token-class 43.6 / 69.9 51.2 / 74.3 57.2 / 72.1 50.7 / 72.1 8.54 / 16.9 40.7 / 45.5 19.4 / 48.2 14.2 / 5.96 20.7 / 29.1
Tü-CL 77.8 / 77.5 74.1 / 80.4 70.0 / 73.4 74.0 / 77.1 11.7 / 21.1 59.9 / 71.8 56.2 / 78.0 85.2 / 85.0 53.3 / 64.0
ByT5unseg 80.8 / 79.7 84.2 / 87.4 78.9 / 82.5 81.3 / 83.2 0.1 / 0.3 42.2 / 53.4 53.7 / 71.0 90.4 / 88.4 46.6 / 53.3
GlossLMunseg, pre 79.8 / 79.2 77.5 / 82.8 76.8 / 80.8 78.0 / 80.9 2.3 / 3.5 1.5 / 1.3 4.1 / 9.6 1.6 / 2.9 2.4 / 4.3
GlossLMunseg, ft 82.1 / 81.5 83.6 / 87.3 78.6 / 81.0 81.4 / 83.3 10.1 / 28.4 57.3 / 64.9 62.8 / 78.9 87.4 / 86.2 56.3 / 65.0
Table 3: Morpheme- and word-level accuracy of various systems on segmented (top) and unsegmented (bottom) text. Best performance per language in each setting the table is bolded. GlossLMall, pre refers to performance using the pretrained GlossLM directly, while GlossLMall, ft refers to performance after fine-tuning the pretrained model on the specific language.
6 Results
Figure 2: Word accuracy for the multilingual GlossLM and various baseline systems on unsegmented data for out-of-domain languages.
Accuracy results for the segmented setting are given in Table 3, and for the unsegmented setting in Table 3 and Figure 2. We report chrF++ scores in Appendix D, finding that they generally show the same trends as accuracy. [^5] All results are for a single run, due to the computational cost of multiple runs.
Segmented setting
In the segmented setting, we observe mixed results. The GlossLM models outperform baselines for certain languages (arp, usp, git) but underperform in other cases (ddo, lez, ntu, nyb). GlossLM matches or outperforms the comparable ByT5 baseline, indicating that the continual gloss pretraining provides small benefits. However, this method does not necessarily overcome other architectures that are more successful, such as the straight-through gradient model Girrbach (2023a) (though future work could utilize our corpus to pretrain models with these architectures). We note that for this setting, performance across models is very high, and it may be difficult to achieve much higher than the observed performances (due to ambiguous morphemes or noisy evaluation data).
Notably, the direct ByT5 model trained on Gitksan far underfits the data, producing improperly formatted data, long random sequences, and outputs including chunks of the prompts. As the Gitksan dataset is very small, the model was unable to learn the correct formatting; meanwhile, the GlossLM model benefits from learning correct formatting from other languages, for a 26.7% morpheme accuracy improvement. For very low-resource documentation projects, our pretrained model shows clear benefits in producing outputs with the correct formatting, while leveraging the multilingual knowledge of ByT5.
We also observe that while finetuning on in-domain languages provides small improvements in performance, the non-finetuned models perform very well. This indicates that multilinguality has not limited performance on each individual language too much, despite the hugely multilingual pretraining corpus. Furthermore, the non-finetuned model is able to make a small percentage of correct predictions (2.9% to 4.9% morpheme accuracy) on unseen languages without training, indicating clear benefits from massively multilingual pretraining.
Unsegmented setting
In the unsegmented setting, the GlossLM models outperform most prior baselines on the in-domain data, although they slightly underperform the ByT5 baseline. This result indicates that with large monolingual datasets, there is not necessarily a benefit to additional multilingual learning on IGT data. This could indicate that the multilingual (non-IGT) corpus used to pretrain ByT5 already sufficiently provides any benefits from crosslingual transfer. The non-finetuned model is also highly competitive, beating prior baselines (except the ByT5 baseline).
For the out-of-domain languages (which tend to have much smaller corpora), the finetuned GlossLM model outperforms the ByT5 baseline on three of four languages and beats previous best baselines by up to 6.6% morpheme accuracy (for Natugu). When the monolingual dataset is very limited, there seems to be a greater benefit from multilingual IGT pretraining, allowing the model to leverage information about segmentation and output formatting from languages in the pretraining corpus. The only case where the finetuned model underperforms SOTA is for Lezgi, slightly underperforming the Tü-CL model. One possibility is that segmentation from surface forms in Lezgi is more difficult, and the hard attention mechanism is beneficial at learning proper morpheme boundaries.
Overall, we suggest the pretrained GlossLM model is a good fit for languages with limited training data, even unseen languages. Directly finetuning a ByT5 model can be very effective for languages with large training corpora, but few documentation projects will have as much available data as the in-domain languages used here (which have 7k to 78k training sentences), and finetuning directly on small datasets risks failure to converge.
Furthermore, we observe that finetuning the GlossLM model on unseen data converges much faster than ByT5–for example, after 50 train epochs for Natugu, the GlossLM-based model achieves 78.1 chrF++ on the eval set, while the ByT5 model is at 30.1. If the desire is to conserve computational resources, using the pretrained model could be beneficial to train efficiently.
7 Error Analysis
7.1 In- vs out-of-vocabulary errors
In-domain Out-of-domain
arp ddo usp git lez ntu nyb
% OOV 30.0 15.6 20.0 78.1 27.3 27.6 8.42
IV 96.2 92.3 91.4 66.7 85.4 91.3 92.8
OOV 55.7 71.7 57.1 26.0 33.9 55.7 32.6
% OOV 30.0 18.7 21.4 80.5 25.5 28.9 9.27
IV 95.3 91.4 89.1 60.0 81.6 91.3 91.7
OOV 50.1 69.7 50.9 20.7 16.4 48.3 30.6
Table 4: Percent of words that are out-of-vocab in the test split for each language along with in- versus out-of-vocabulary accuracy at the word level. Top is the segmented setting (GlossLMall, ft), bottom is unsegmented (GlossLMunseg, ft).
We are interested in how well our models can generalize to unseen words, which are often novel combinations of new and previously-seen morphemes. We report word-level accuracy for our finetuned GlossLM models, indicating whether the transcribed word is in- or out-of-vocabulary in Table 4, as well as the percent of OOV words in the test set. The OOV rate between segmented and unsegmented may vary slightly, as mappings between segmented and unsegmented forms are not necessarily one-to-one. We consider a word to be in-vocabulary if the form of the word in the transcription and its corresponding gold label in the gloss co-occur in the training data. We also include morpheme-OOV rates and statistics on lexical gloss overlap with translations in Appendix E.
Across most languages, our models perform well on in-vocabulary words (with Gitksan being a notable exception, likely due to the small amount of training data). Within the OOV split, we expect that models perform better on languages that are agglutinative with lower morpheme-OOV rates. We find that this is the case, as Tsez (which has the highest OOV accuracy) is a highly agglutinative language with the lowest rate of OOV morphemes in the test set, implying that the majority of OOV words are actually composed of seen morphemes. However, Lezgi—which is also highly agglutinative—has strikingly low OOV accuracy; this may be due to the highly complex inflections for case marking, which could be more difficult for our system to learn. Finally, Nyangbo also shows relatively low OOV word accuracy, possibly due to the lack of translations in the input, which could help predict OOV stem labels.
7.2 Qualitative Analysis of Errors
A manual analysis of GlossLM outputs reveal two interesting points of consideration.
In certain cases, the model may rely too heavily on translations when predicting lexical glosses.
When inspecting the Uspanteko outputs from the finetuned GlossLMall model, we found that the model outputs “esta@bien” (the correct label for the morpheme tziin) as the label for jaan, which should be glossed as “bueno”; this happens especially often when the phrase “esta bien” appears in the translation. In other cases, the model sometimes outputs a word or phrase present in the translation, even if there is no occurrence of that item in the gold gloss. We include example system outputs to illustrate these cases in Appendix F.
Our metrics potentially underestimate GlossLM’s performance on the glossing task.
There are sometimes inconsistencies in the IGT labels where multiple interchangeable glosses are used for the same morpheme, particularly for lexical glosses (e.g. the Arapaho word ’eeneisih’i is glossed in the data as “how.X.things.are.named”, “how.s.o..is.named”, and “how.things.are.named”). These types of errors can be accounted for when using chrF++, but another issue that may more greatly impact generative models—especially if it learns to use translations effectively—is if the model outputs a semantically similar label that is dissimilar in form; for example, outputting “PREP” (preposition label) instead of “en” (“in” in Spanish, which is a preposition). We include examples of such outputs in Appendix F.
8 Related Works
8.1 Automatic Interlinear Glossing
Recent research has explored various methods for generating IGT, including rule-based methods (Bender et al., 2014), active learning (Palmer et al., 2010, 2009), conditional random fields (Moeller and Hulden, 2018; McMillan-Major, 2020), and neural models (Moeller and Hulden, 2018; Zhao et al., 2020). The 2023 SIGMORPHON Shared Task (Ginn et al., 2023) compared a number of highly-effective IGT generation systems, including ensembled LSTMs (Coates, 2023), straight-through gradient estimation (Girrbach, 2023a), CRF-neural systems (Okabe and Yvon, 2023a), and BiLSTM encoders (Cross et al., 2023).
In particular, this work is inspired by He et al. (2023), which pretrains ByT5 models on the ODIN corpus, and Okabe and Yvon (2023b), which pretrains a CRF model on the IMTVault corpus. However, neither explore using a pretraining corpus as large as ours, nor do they evaluate on unsegmented text. Furthermore, neither of these studies find significant benefits to using pretraining corpora, while we observe benefits under certain conditions.
8.2 Large Multilingual Pretrained Models
Prior work has shown that large, massively multilingual pretrained language models can boost performance across low- and high-resource languages on a variety of tasks. These include encoder-decoder models trained with the masked language modeling objective Pires et al. (2019); Conneau et al. (2020) and the span corruption objective Xue et al. (2021, 2022), as well as decoder-only language models Workshop et al. (2023); Shliazhko et al. (2024). Work such as Wang et al. (2020), Adelani et al. (2022b), and Adelani et al. (2022a) has shown that continual pretraining and/or finetuning large multilingual models is an effective method for tasks like low-resource language translation and named entity recognition.
9 Conclusion
Automatic generation of interlinear glossed text has great potential to aid documentation projects, but limited monolingual data limits effectiveness. In this research, we attempt to overcome these limitations using crosslingual transfer, compiling the largest multilingual IGT dataset and pretraining massively multilingual glossing models. Our models outperform any existing work in the majority of settings, demonstrating clear benefits of multilingual representations–particularly for languages with very small corpora. We make our dataset and models available for future research, in hopes that they will aid future NLP research and benefit documentation projects.
10 Limitations
Our work evaluates the effectiveness of massively multilingual pretraining on seven IGT datasets in different languages using the ByT5 architecture. We did not experiment with pretraining on other architectures, which may show different results. While we believe the selected evaluation languages cover a diverse set of features and dataset sizes, other languages may show better or worse results.
Our pretraining corpus consists of all IGT data we were able to find and utilize. As such, it is not evenly distributed among languages, over-representing a few languages with large language documentation efforts. Thus, models pretrained on the corpus will perform better on these and similar languages.
The only hyperparameter optimization we performed was finding a batch size that fit our GPUs and tuning epochs and early stopping in order to ensure convergence. We did not conduct hyperparameter search over other parameters such as learning rate or optimizer, architecture parameters, or dataset splits.
When evaluating predictions, we ignored punctuation (as our primary concern was gloss accuracy). Certain models may perform better or worse at outputting proper punctuation format, which could be a concern for certain applications.
Finally, it has been demonstrated that IGT generation models are often not robust to domain shift, compared with human annotators (Ginn and Palmer, 2023). Our models will likely have impacted performance for out-of-domain texts, such as highly technical or domain-specific language.
11 Ethical Considerations
We hope this work can aid in the struggle against language extinction. However, language documentation, preservation, and revitalization require far more than generating IGT, and we should be careful not to understate the difficulty of these efforts. We utilize datasets produced by the painstaking effort of language documenters and speakers, and strive to treat the corpora as human artifacts, not just data to be consumed.
We hope our research can aid documentary linguists through automated gloss prediction. However, we caution against using these systems without human collaboration, as they can introduce error and miss novel linguistic insights. There is some risk of these systems being used to replace human annotators, which we strongly oppose.
While we try to train only the necessary models for our experiments, training large machine learning models carries an environmental cost (Bender et al., 2021; Strubell et al., 2020).
We do not evaluate our corpus for bias (racial, gender, etc) or inclusive language, and it’s possible that our models can carry some of these biases.
Finally, NLP work that involves Indigenous and endangered languages has historically been plagued by colonialist approaches to data use and technology development (Schwartz, 2022). The large IGT datasets for endangered languages (Arapaho, Guarani, Uspanteko) were collected in collaboration with native communities, and our work is in accordance with the agreements for their usage.
Appendix A Language Distribution
Figure 3: Counts per language. We only show languages with at least 2k samples present in the dataset. Arapaho (arap1274) is by far the most represented language in our data, followed by Uspanteko (uspa1245). Both languages are part of the SIGMORPHON Shared Task dataset.
Figure 4: Counts per language family in our dataset. We only show language families with at least 1k samples present in the dataset. Language isolates and languages without recorded families in Glottolog Hammarström et al. (2023) are categorized under “None”.
In Figure 3 and Figure 4, we visualize the number of examples per language and language family on a portion of our dataset.
Appendix B Grambank Typological Analysis
Along with number of languages, we would also like to measure whether the distribution of typological features in our dataset is reflective of the diversity of features in the world. We use the Grambank typological database Skirgård et al. (2023) as a standard against which to judge the typological diversity of our dataset. Grambank covers over 2430 languages, with up to 195 features (e.g., "What is the order of numeral and noun in the NP?") per language. The values of the features comprise a vector for each language.
43% of our languages are found in Grambank, amounting to 72% coverage over all training instances. However, Grambank does not have complete feature vectors for all languages. Using the method described by Skirgård et al. (2023), we imputed missing feature values (9.7% of all features), resulting in complete feature vectors of size 161, as we only accept features that are defined for at least $64\%$ of our dataset (as to balance dataset coverage as well as feature coverage).
We then create an average feature vector for our dataset by averaging the feature vectors of the the languages present in Grambank (weighting the average based on the number of training instances in each language) and compare this to the average of feature vectors for all languages in Grambank. We find a cosine similarity of 0.92 between the two vectors. In comparison, the language with the greatest similarity to the average Grambank vector in our data has a cosine similarity of 0.81. We report additional details of our methods and analysis below.
B.1 Imputation Details
We adapt the imputation procedure described in Skirgård et al. (2023) and follow the steps below. Thresholds were chosen to maximize the language coverage while keeping the imputed values below 10%.
  • Removed languages that had > 36% missing data out of the dataset
  • Removed features that had >36% missing data among the remaining languages
  • Binarized the multistate values
  • Removed all but one dialect for each language according
  • Imputed missing values with iterated random forest with MissForest [^6] https://rpubs.com/lmorgan95/MissForest
B.2 Underrepresented Features
ID Feature Average Value
GB024b Does the numeral always follow the noun 0.468
in the NP?
GB193b Are most adnominal property words placed 0.536
after nouns?
GB025b In the pragmatically unmarked order, does the 0.425
adnominal demonstrative follow the noun?
GB118 Are there serial verb constructions? 0.406
GB130a Is the order in intransitive clauses with 0.382
a full nominal subject consistently SV?
Table 5: Most underrepresented Grambank features in training data with their average value (from imputed vectors).
In Table 5 we report the top five features for which the average representation in our training data is most distant from the expected value (as determined by averaging feature values across all languages in Grambank).
B.3 Averaged Feature Vector
For reference, we include the vector set of 161 average Grambank features over all training languages weighted by the number in Table 11.
Appendix C Training Hyperparameters
Training the GlossLMall and GlossLMunseg models used A6000 and A100 GPUs, and took around 5 days per run. We list the hyperparameters used in Table 6.
Parameter Value
Optimizer Adafactor
Initial LR 5e-5
Weight decay 0.01
Batch size 2
Gradient accumulation steps 64
Epochs 13
Table 6: Pretraining Hyperparameters
Appendix D chrF++ Scores
We provide chrF++ scores for various systems in Table 10. Broadly, we found the results supported the same trends as the accuracy measurements, although the magnitude of the differences varied.
Appendix E Test Split Vocabulary Statistics
Language Morpheme %OOV
arp 3.6
ddo 41.2
usp 4.9
git 2.8
lez 1.1
ntu 0.5
nyb 5.3
Table 7: Percent of out-of-vocabulary morphemes in the test split.
Language OOV Token Recall
arp 49.87
ddo 44.13
usp 44.89
git 58.21
lez 71.66
ntu 40.14
Table 8: Percent of lexical glosses present in the translation in the test split. Nyangbo examples do not include translations.
We include additional statistics for morpheme OOV rate and recall rate for lexical labels in the translation in Table 7 and Table 8, as reported in Ginn et al. (2023).
Appendix F Example System Outputs
Sample ID
uspa1245_136 transcription jaan
translation Esta bien.
gold bueno
output esta@bien
uspa1245_149 transcription kond (ti’) laj chaak
translation Cuando en su trabajo.
gold cuando ??? PREP trabajo
output cuándo??? PREP trabajo
lezg1247_1 transcription вич- ни хьун- нва- й къвалах- ар я
translation and all this were real stories
gold reflxv- FOC - be- PERF- PST- word- PL was
output himself- FOC be- PERF- PTP real.story- PL was
lezg1247_1 transcription ва гьада гьикъван гьа терези- ди- н стрелка пара хкаж хьа- нва- тІа гьам вилик кутуна- нва
translation she put first that person which is more valuable according to the position of scale arrow.
gold and then how that.the.same scale- DIR- GEN arrow very up happened- PERF- COND that before put- PERF
output that according.to that.the.same value- OBL- GEN position.in very hit be- PERF- COND that.the.same behind put- PERF
arap1274_991 transcription neetotohoe
translation Take off your pants !
gold take.off.one’s.pants
output take.off.pants
arap1274_1998 transcription cee’iyo
translation payday .
gold payment
output pay.day
arap1274_1667 transcription howouunonetiit hiniito’eibetiit biisiinowoot niihooku’oot beh- ’entou- ’
translation pity , relationships , learning through observation , watching closely , it’s all there .
gold pity.mercy relatedness learning.by.observing watching.along all- located.present 0S
output pity.mercy relationships learning.through.observation watching.along all- located.present- 0S
Table 9: Selected example outputs to illustrate errors by GlossLMall finetuned models.
We include examples of system outputs to illustrate the types of errors discussion in subsection 7.2 in Table 9.
Model chrF++ Score (Segmented)
In-domain Out-of-domain
arp ddo usp Avg git lez ntu nyb Avg
Top-choice 75.0 71.9 71.4 72.8 33.7 69.7 74.5 62.2 60.0
Token-class 84.2 81.8 75.3 80.4 25.6 52.1 65.4 84.3 56.9
Tü-CL 85.2 88.4 77.7 83.8 34.8 78.9 87.9 87.7 72.3
CRF 84.2 88.2 79.4 83.9 40.9 79.2 88.4 84.3 73.2
SMF 80.7 86.6 76.3 81.2 28.0 62.0 78.8 82.2 62.6
ByT5all 84.2 91.8 84.6 86.9 9.07 74.9 85.1 88.8 64.5
GlossLMall, pre 85.2 90.8 83.1 86.4 21.5 14.1 18.2 8.8 15.7
GlossLMall, ft 86.3 91.5 85.9 87.9 43.1 75.0 84.1 87.4 72.4
Model chrF++ Score (Unsegmented)
In-domain Out-of-domain
arp ddo usp Avg git lez ntu nyb Avg
Top-choice 44.0 63.5 55.1 54.2 8.4 51.6 40.5 74.0 43.6
Token-class 56.2 72.9 65.3 64.8 18.8 56.4 45.1 18.8 34.8
Tü-CL 77.6 84.6 72.5 78.2 23.0 71.5 78.6 84.1 64.3
ByT5unseg 80.7 90.0 83.0 84.6 7.6 59.6 77.0 88.4 58.2
GlossLMunseg, pre 80.5 86.8 81.0 82.8 19.4 13.3 16.3 8.1 14.3
GlossLMunseg, ft 82.9 89.8 81.7 84.8 34.9 68.8 80.7 85.5 67.5
Table 10: chrF++ scores of various systems on segmented (top) and unsegmented (bottom) data. Best performance per language in each setting the table is bolded. GlossLMall, pre refers to performance using the pretrained GlossLM directly, while GlossLMall, ft refers to performance after finetuning on the specific language.
ID Avg. Value ID (cont.) Avg. Value (cont.) ID (cont.) Avg. Value (cont.)
GB020 0.183 GB111 0.573 GB305 0.364
GB021 0.15 GB113 0.6 GB309 0.185
GB022 0.199 GB114 0.502 GB312 0.649
GB023 0.058 GB115 0.52 GB313 0.161
GB026 0.065 GB116 0.007 GB314 0.038
GB027 0.351 GB117 0.705 GB315 0.054
GB028 0.373 GB118 0.123 GB316 0.031
GB030 0.21 GB119 0.196 GB317 0.03
GB031 0.067 GB120 0.276 GB318 0.1
GB035 0.484 GB121 0.215 GB319 0
GB036 0.023 GB122 0.162 GB320 0.002
GB037 0.019 GB124 0.081 GB321 0.065
GB038 0.036 GB126 0.369 GB324 0.07
GB039 0.165 GB129 0.001 GB326 0.487
GB041 0.044 GB131 0.329 GB327 0.563
GB042 0.097 GB132 0.519 GB328 0.515
GB043 0.021 GB133 0.25 GB333 0.721
GB044 0.602 GB134 0.704 GB334 0.276
GB047 0.701 GB135 0.729 GB335 0.093
GB048 0.634 GB136 0.235 GB336 0.002
GB049 0.661 GB137 0.206 GB408 0.352
GB051 0.18 GB138 0.43 GB409 0.073
GB052 0.02 GB139 0.653 GB410 0.532
GB053 0.391 GB140 0.304 GB415 0.233
GB054 0.02 GB147 0.567 GB430 0.013
GB057 0.167 GB148 0.05 GB431 0.331
GB058 0.019 GB149 0.299 GB432 0.263
GB059 0.134 GB151 0.018 GB433 0.389
GB068 0.484 GB152 0.138 GB519 0.168
GB069 0.392 GB155 0.578 GB520 0.094
GB070 0.269 GB156 0.032 GB521 0.05
GB071 0.326 GB158 0.526 GB024a 0.664
GB072 0.538 GB159 0.176 GB024b 0.132
GB073 0.296 GB160 0.6 GB025a 0.678
GB074 0.543 GB165 0 GB025b 0.169
GB075 0.556 GB166 0.004 GB065a 0.644
GB079 0.51 GB167 0.036 GB065b 0.184
GB080 0.69 GB170 0.452 GB130a 0.469
GB081 0.032 GB171 0.179 GB130b 0.332
GB082 0.544 GB172 0.086 GB193a 0.626
GB083 0.647 GB177 0.272 GB193b 0.202
GB084 0.498 GB184 0.507
GB086 0.653 GB185 0.292
GB089 0.523 GB186 0.101
GB090 0.146 GB192 0.098
GB091 0.539 GB196 0.054
GB092 0.106 GB197 0.041
GB093 0.394 GB198 0.409
GB094 0.143 GB257 0.519
GB095 0.054 GB260 0.097
GB096 0.018 GB262 0.095
GB098 0.369 GB263 0.181
GB099 0.113 GB264 0.062
GB103 0.368 GB285 0.009
GB104 0.286 GB286 0.154
GB105 0.517 GB291 0.013
GB107 0.521 GB297 0.059
GB108 0.414 GB298 0.083
GB109 0.107 GB299 0.314
GB110 0.163 GB302 0.131
Table 11: Grambank Feature Averages over Training Set

### v2

@@
-GlossLM: Multilingual Pretraining for Low-Resource Interlinear Glossing
+GlossLM: A Massively Multilingual Corpus and Pretrained Model for Interlinear Glossed Text
 Abstract
-A key aspect of language documentation is the creation of annotated text in a format such as interlinear glossed text (IGT), which captures fine-grained morphosyntactic analyses in a morpheme-by-morpheme format.
-Prior work has explored methods to automatically generate IGT in order to reduce the time cost of language analysis.
-However, many languages (particularly those requiring preservation) lack sufficient IGT data to train effective models, and crosslingual transfer has been proposed as a method to overcome this limitation.
+Language documentation projects often involve the creation of annotated text in a format such as interlinear glossed text (IGT), which captures fine-grained morphosyntactic analyses in a morpheme-by-morpheme format.
+However, there are few existing resources providing large amounts of standardized, easily accessible IGT data, limiting their applicability to linguistic research, and making it difficult to use such data in NLP modeling.
 We compile the largest existing corpus of IGT data from a variety of sources, covering over 450k examples across 1.8k languages, to enable research on crosslingual transfer and IGT generation.
-Then, we pretrain a large multilingual model on a portion of this corpus, and further finetune it to specific languages.
-Our model is competitive with state-of-the-art methods for segmented data and large monolingual datasets.
-Meanwhile, our model outperforms SOTA models on unsegmented text and small corpora by up to 6.6% morpheme accuracy, demonstrating the effectiveness of crosslingual transfer for low-resource languages.111 Our pretrained model and dataset are available at https://huggingface.co/lecslab/glosslm and https://huggingface.co/datasets/lecslab/glosslm-corpus-split, and will be made accessible through a web interface for use in language documentation efforts.
+We normalize much of our data to follow a standard set of labels across languages.
+Furthermore, we explore the task of automatically generating IGT in order to aid documentation projects.
+As many languages lack sufficient monolingual data, we pretrain a large multilingual model on our corpus.
+We demonstrate the utility of this model by finetuning it on monolingual corpora, outperforming SOTA models by up to 6.6%.
+We will make our pretrained model and dataset available through Hugging Face, as well as provide access through a web interface for use in language documentation efforts.
 […]
 […]
-GlossLM: Multilingual Pretraining for Low-Resource Interlinear Glossing
-[…]
+GlossLM: A Massively Multilingual Corpus and Pretrained Model for Interlinear Glossed Text
+Michael Ginn*1  Lindia Tjuatja*2  Taiqi He2  Enora Rice1 Graham Neubig2Alexis Palmer1Lori Levin2 1University of Colorado Boulder  2Carnegie Mellon University michael.ginn@colorado.edulindiat@andrew.cmu.edu * Equal contribution
 1 Introduction
-footnotetext: * Equal contribution
 With nearly half of the world’s 7,000 languages considered endangered, communities of minoritized language speakers are working to preserve and revitalize their languages (Seifart et al., 2018).
-These efforts often involve the collection, analysis, and annotation of linguistic data.
-Annotated text is especially useful in the creation of reference texts (like dictionaries and grammars) and other archival materials, as well as in the development of language technologies including searchable digital text Blokland et al.
+These efforts often involve collection, analysis, and annotation of linguistic data.
+Annotated text can be used in the creation of reference materials (such as dictionaries and grammars) as well as to develop language technologies including searchable digital text Blokland et al.
 (2019); Rijhwani et al.
@@
 (2017); Chaudhary et al.
-(2023)
+(2023).
 Figure 1: Components of interlinear gloss with an Arapaho sentence and English translation Cowell (2020).
@@
 The task of automatic glossing uses some or all of the information in the gray box (transcription & translation) to generate the gloss line.
-One prevalent form of linguistic annotation in language documentation projects is interlinear glossed text (IGT).
-IGT is a multi-line data format which includes (1) a transcription of speech in the language, (2) an aligned morpheme-by-morpheme description, and oftentimes (3) a free translation.
-An example of each of these components is illustrated in Figure 1.
-IGT is useful for illustrating morphosyntactic features of languages that other linguists or researchers may not be familiar with, and can also be used as examples for grammatical sketches or language learning textbooks, among other applications.
-Creating IGT corpora is time-consuming and difficult, requiring documenters to perform linguistic analysis and extensive documentation simultaneously, often for languages with limited prior documentation.
-Research has found that computational tools can help accelerate the annotation process and overcome this bottleneck Palmer et al.
-(2009).
-In particular, models that take as input transcriptions and translations (which don’t require linguistic training to produce) and output glosses can rapidly bootstrap creation of complete IGT datasets.
-The majority of prior work on automatic glossing focuses on training monolingual models that can predict IGT for a single language (Moeller and Hulden, 2018; McMillan-Major, 2020; Zhao et al., 2020).
-This approach has two major drawbacks.
-First, many language community members or linguists do not have the expertise, time, or resources to create special-purpose models, creating a high barrier of entry for adoption.
-Second, in the rare cases where expertise is available, there is often insufficient amounts of data to train effective monolingual models, with many IGT corpora having only a few thousand examples or less.
-In this work, we aim to overcome the monolingual data bottleneck by creating a multilingual pretrained gloss model that can be further adapted to specific languages and gloss formats.
-Inspired by recent advances in massively multilingual language modeling, we compile various existing IGT corpora into a dataset containing over 450k IGT examples in 1.8k languages, making it the largest existing collection of digitized IGT.
-We then train our model, GlossLM, by continually pretraining ByT5 Xue et al.
-(2022), and evaluate it on seven typologically diverse languages.
-We find that our model is highly competitive with existing models, outperforming them broadly on the more challenging task of glossing unsegmented text (where words are not split into individual morphemes), as well as in certain cases where segmented text is provided.
-Furthermore, we observe that our model retains high performance on seen languages even without finetuning, indicating that it can be used as-is without specialization to particular languages.
+Interlinear glossed text (IGT) is a widespread format in language documentation for linguistic annotation.
+IGT is a multi-line data format (see Figure 1) which includes (1) a transcription of speech in the language, (2) an aligned morpheme-by-morpheme description, and oftentimes (3) a free translation.
+IGT can be used to illustrate morphosyntactic features of languages that other researchers may not be familiar with, and it is a popular format for examples in linguistics papers and textbooks.
+It also serves as a resource in the NLP context for the creation of morphological paradigms (Moeller et al., 2020), machine translation (Zhou et al., 2019), generating precision grammars (Bender et al., 2013), and other tools including POS taggers and dependency parsers (Georgi, 2016).
+Compiling IGT Data
+Though IGT often follows a common glossing format, gloss conventions vary wildly.
+Furthermore, IGT data is rarely compiled into large, standardized corpora, often existing as scattered examples in research papers.
+To address this, we compile the largest corpus of digitized IGT from various existing sources, with over 450k examples in 1.8k languages (§3).
+We explore methods to normalize this data (§4)., standardizing over 80% of the grammatical glosses in the corpus to follow the UniMorph (Sylak-Glassman, 2016).
+We will release our corpora for future NLP, linguistics, and language documentation research.
+Automating IGT Generation
+Futhermore, creating new IGT corpora is difficult and time-consuming, requiring documenters to perform linguistic analysis and extensive documentation simultaneously.
+Research has found that computational tools can help accelerate annotation and overcome this bottleneck Palmer et al.
+(2009) by predicting the gloss line of IGT given a transcription.
+The majority of prior work on automatic glossing focuses on training monolingual models that can predict IGT for a single language (Moeller and Hulden, 2018; McMillan-Major, 2020; Zhao et al., 2020), however, these models can struggle when data is limited and require dedicated effort to train and deploy.
+We aim to overcome the monolingual data bottleneck by creating a multilingual pretrained glossing model that can be adapted to specific languages and gloss formats.
+We continually pretrain a model on our corpus, finding that pretrained multilingual model retains high performance across languages.
+We then finetune the pretrained model on monolingual data, including languages that do not appear in the pretraining corpus.
+We achieve a new SOTA performance on five out of seven languages, and demonstrate clear improvements for low-resource language settings over an equivalent finetuned model without our continual pretraining (§7).
 2 Interlinear Glossed Text (IGT)
-2.1 Description of IGT Format
-IGT is a structured data format which presents a transcription of text in a language being studied, along with its gloss—aligned labels that mark each morpheme’s meaning and/or grammatical function—in a different language.
-Oftentimes, a free translation in a widely-spoken language is included.
-An IGT example for Arapaho is given in subsection 2.1 Cowell (2020), with glosses and translations in English:
+2.1 Format
+Interlinear glossed text is a structured data format which presents text in a language being studied along with morphological glosses—aligned labels that indicate each morpheme’s meaning and/or grammatical function.
+Often, a free translation in a widely-spoken language is included as well.
+An IGT example for Arapaho is given in subsection 2.1 Cowell (2020), with glosses and translations in English.
 […]
@@
 Each morpheme in the Arapaho sentence (e.g.
-tih) is directly aligned with glosses (separated by periods) that describe that morpheme’s function and/or meaning.
+tih) is directly aligned with a gloss (e.g.
+when.Past) that describes the morpheme’s function and/or meaning.
 Labels in all caps (e.g.
@@
 speak) are lexical glosses.
-IGT examples may also use unsegmented transcriptions, as in the Uspanteko example in subsection 2.1 Pixabaj et al.
-(2007).
-In this case, words and their labels are aligned, but no explicit alignment between morpheme labels and the original morphemes themselves is present, and thus the segmentation of words into morphemes is unclear:
+Periods are used for fusional morphemes, which carry several morphological or lexical functions in a single morpheme.
+IGT examples may instead use unsegmented transcriptions, as in the Uspanteko example in subsection 2.1 Pixabaj et al. (2007).
 […]
@@
 […]
-IGT is highly valuable to documentation projects, and has also been used in the NLP context for the creation of morphological paradigms (Moeller et al., 2020), machine translation (Zhou et al., 2019), generating precision grammars (Bender et al., 2013), and other tools including POS taggers and dependency parsers (Georgi, 2016).
-2.2 Challenges with Automatic Glossing
-Effective glossing requires knowledge of the target language as well as linguistic understanding of morphological patterns.
+Here, words and their labels are aligned, but no explicit alignment between morpheme glosses and individual morphemes is provided, and thus the segmentation of words into morphemes is unclear.
+2.2 Challenges with Interlinear Glossing
+Effective glossing requires expert knowledge of the target language and linguistic understanding of morphological patterns.
 Furthermore, certain factors exist that make this task particularly difficult for automated systems.
-Lack of segmented transcriptions: As in subsection 2.1, often only an unsegmented transcription is available.
-In these cases, an automated system cannot predict glosses morpheme-by-morpheme, but must instead learn to perform simultaneous segmentation and glossing.
-Variance in glossing styles: While there are widely-used standards for glossing conventions, such as the Leipzig glossing rules Comrie et al.
-(2008), exact glossing styles may vary widely from documenter to documenter Chelliah et al.
-(2021).
-Documentation projects may have differences in label spelling (e.g.
-SING/SG/S to denote singular), formatting (e.g.
-underscores/periods for multi-word lexical glosses), and language specific labels (e.g.
-a linguist may want to specify different levels of an intensifier marker), among others.
-Extremely low-data settings: When IGT is used for language documentation, the target languages of interest are likely those with very small amounts of digital data to begin with.
-Given these challenges, we want our model to be easily adaptable to different languages with varying amounts of training data, as well as able to handle both segmented and unsegmented inputs.
-3 Dataset
+Often, transcriptions are not segmented into morphemes, and systems must perform simultaneous segmentation and glossing.
+Glossing conventions and formats may vary widely from documenter to documenter Chelliah et al.
+(2021), with differences in label spelling (e.g.
+Sing/Sg/S to denote singular), formatting and punctuation, and language-specific labels.
+Furthermore, nearly all languages have very little digitized IGT data, posing difficulty to automated systems and human annotators alike.
+Finally, even when automated systems have been created, practical deployment remains an additional challenge for documenters.
+3 GlossLM Corpus
 We compile and clean the largest IGT dataset from a variety of sources and languages.
@@
 Guarani Corpus 1 803
+Total 1,782 250,582
 Table 1: Number of unique examples and languages in each source corpus for the GlossLM dataset.
@@
 We use the preprocessed version of ODIN by He et al.
-(2023), which discards languages with less than five IGT samples, resulting in 84k unique glossed sentences across 936 languages.
+(2023), which discards languages with fewer than five IGT samples, resulting in 84k unique glossed sentences across 936 languages.
 SIGMORPHON Shared Task
@@
 In total we have 250k unique IGT instances in 1.8k languages.
-We use Hugging Face Datasets to store our corpus in a widely-used, publicly accessible format.
-Each row retains a unique ID and reference to its source, so examples can be easily cross-referenced.
 If datasets explicitly indicate whether an IGT line is segmented, we use this value.
@@
 (2023) in Appendix B.
+4 Normalizing Gloss Labels
+4.1 Motivation
+As our data comes from a variety of sources, spanning many languages and documentation projects, there is a great amount of diversity of the morphological glosses used.
+This includes cases where several different labels are used to indicate the same feature (e.g.
+Sing, Sg, or S for singular), as well as formatting differences such as the usage of periods (e.g.
+1Sg vs 1.Sg).
+We explore the value and feasibility of normalizing glosses to a single standardized format.
+On one hand, normalizing glosses may make it easier to train models that utilize crosslingual information through shared gloss labels, but on the other hand, it can be difficult to select a single schema that preserves the original intent of all annotators.
+We split gloss lines by period and count the number and frequency of unique grammatical gloss labels across our corpus (focusing on the all-caps functional glosses, not stems) and visualize the distribution in Figure 2.
+Figure 2: Distribution of unique glosses across all languages.
+There are 11,493 unique glosses which roughly form a Zipfian distribution (Zipf, 1945).
+The most common glosses, unsurprisingly, are labels such as Pl (plural, 52,488 instances), 3Sg (3rd-person singular, 39,147), and Past (36,124), which occur broadly across many languages.
+Normalizing all unique glosses would be a monumental task with uncertain benefits.
+However, we observe that the 200 most common gloss types account for 82.7% of glosses in our dataset.
+We focus on normalizing these glosses: e.g.
+all instances of Past and Pst, which are both in the top 200, should be normalized to the same label.
+We note that there are other aspects of the data that vary (e.g.
+periods vs.
+underscores for multi-word glosses, representing non-concatenative morphology), as addressed in Mortensen et al.
+(2023).
+Future work could potentially focus on the benefits of these aspects of normalization on training.
+4.2 Methodology
+We select the UniMorph schema of Sylak-Glassman (2016) as our standardized set.
+While no single set of labels captures the intricacies of all of the world’s languages, UniMorph is widely used and has coverage for many common features.
+The two lead authors of this paper jointly created a mapping from the labels in our dataset to UniMorph labels.
+While many mappings were obvious (or already compatible with UniMorph), others posed a myriad of issues.
+For glosses primarily used for a single language, we consulted the original source dataset to determine the meaning.
+Ambiguous labels
+Several labels were ambiguous, corresponding to one of several UniMorph glosses depending on the language and annotator.
+For example, the label S (appearing 20,855 times) is used for singular, subject, and noun/sustantivo (at least) in our dataset.
+In order to map these, we would need to analyze on a sentence-by-sentence basis, which was not feasible; thus, we left these labels as-is.
+Glosses not in UniMorph
+UniMorph primarily focuses on common crosslingual inflectional features, and does not cover the full extent of the morphological systems of the world.
+We observed 64 of 200 (32%) gloss labels with no clear UniMorph equivalent, including demonstratives (Dem, 15,585 instances), obliques (Obl, 13,639), and clitics (Cl, 7,453).
+In many cases, there is a related UniMorph gloss that is more general or more specific; for example preterite (Pret, 1,986) could be mapped to simple past (Pst).
+However, this would be an imprecise mapping, and could be confusing to a linguist of the particular language, so we again elect to leave these glosses unmapped.
+We use our mapping to normalize the dataset and will make the normalized version available in addition to the original.
+4.3 Use in Future Research
+We believe our dataset can potentially be useful across NLP research, linguistic research, and language documentation.
+NLP researchers benefit from a single, easily-accessible dataset covering many languages, which can be used for future research on interlinear glossing, morpheme analysis, segmentation, and translation.
+Linguistics researchers will be able to use the dataset to easily search for phenomena across languages, particular with the normalized version of the dataset.
+For example, a linguist could find examples of sentences demonstrating the ergative/absolutive distinction.
+They could further refine this analysis by narrowing the results to related languages, using the glottocodes in our dataset.
+As another example, if a linguist wished to determine how prior researchers have annotated examples in a particular language, they would previously have to search across research papers, textbooks, and small corpora.
+With our dataset, it is trivial to pull up all of the examples in a particular language, potentially compiled from many sources.
+5 Automatic IGT Generation
+Next, we evaluate the applicability of our dataset to the NLP task of automatic gloss prediction.
+We select the IGT data from the SIGMORPHON Shared Task on Interlinear Glossing (Ginn et al., 2023) to use for evaluation and testing, as this data consistently adheres to a set of glossing conventions and has been evaluated on prior models.
 5.1 Target Languages
-We select data from the SIGMORPHON Shared Task to use for evaluation and testing, as this data is highly consistent and has been evaluated on prior models.
-We use the same splits as the shared task (Table 2) and designate three languagues–Arapaho, Tsez, and Uspanteko–as in-domain languages, which are included in the pretraining corpus.
-We use Gitksan, Lezgi, Natugu, and Nyangbo as out-of-domain languages, which do not occur in the pretraining corpus, but are used to finetune models.
+We reuse the train/eval/test splits for the seven languages from the SIGMORPHON Shared Task (Table 2).
+We designate three languagues—Arapaho, Tsez, and Uspanteko—as in-domain languages, which are included in the GlossLM corpus.
+We use Gitksan, Lezgi, Natugu, and Nyangbo as out-of-domain languages, which are omitted from the corpus.
 All languages except Nyangbo include translations.
+The shared task included two distinct settings:
+  • In the open track, the transcription lines were segmented into morphemes.
+This becomes a token classification task, and tends to be far easier, with SOTA models achieving 80-90% accuracy (and even a naïve method that simply selects the most common gloss for each morpheme was very effective).
+  • In the closed track, transcription lines were not segmented.
+This setting is far more challenging, as models must jointly learn to segment words and predict glosses, and the best models achieved as low as 11% accuracy on small datasets.
+On the other hand, these models have the potential to be more valuable to documentation projects, where segmented text may not be available.
+In our experiments, we focus on the unsegmented setting (closed track). However, the segmented data is also included in the GlossLM corpus, and could easily be used in future research.
 Finetuning
 Language Pre-train Train Eval Test
-Other languages 313790 - - -
+Other languages 313,790 - - -
 In-domain languages
-Arapaho (arp) 78,264 78,264 9,784 9,784
-Tsez (ddo) 7116 7116 890 890
-Uspanteko (usp) 19,584 19,584 464 1,266
+Arapaho (arp) 39,132 39,132 4,892 4,892
+Tsez (ddo) 3,558 3,558 445 445
+Uspanteko (usp) 9,774 9,774 232 633
 Out-of-domain languages
-Gitksan (git) - 163 84 74
-Lezgi (lez) - 1,411 176 174
-Natugu (ntu) - 1,582 198 198
-Nyangbo (nyb) - 4,200 526 526
-Table 2: Number of total (segmented and unsegmented) pretrain (for in-domain languages), train, evaluation, and test examples for the target languages.
+Gitksan (git) - 74 42 37
+Lezgi (lez) - 705 88 87
+Natugu (ntu) - 791 99 99
+Nyangbo (nyb) - 2,100 263 263
+Table 2: Number of total (unsegmented) pretrain (for in-domain languages), train, evaluation, and test examples for the target languages.
+5.2 Evaluation Metrics
+For evaluating predictions, we strip punctuation (except for within glosses).
+[^2] Because of this post-processing, our results for baselines are slightly different than what original sources report.
+We evaluate morpheme accuracy, which counts the number of correct morpheme glosses in the correct position.
+Hence, if a gloss is incorrectly inserted or deleted, the subsequent glosses will be incorrect.
+We also evaluate word accuracy, which counts the number of entire correct word glosses.
+However, accuracy may sometimes be too strict of a measurement—especially for generative models—as minor character insertions/deletions in the label are penalized heavily.
+Thus, we also evaluate chrF++ (Popović, 2015), a character-level metric often used in machine translation.
+chrF++ measures the F1 score over character n-grams between the reference and predictions, and is robust to insertions and deletions, unlike accuracy.
 6 GlossLM Model
-Using our IGT dataset, we aimed to create a single multilingual pretrained model that could be easily adapted to documentation projects, both for known languages and unseen ones.
+Using our IGT corpus described in section 3, we train a single multilingual pretrained model for the glossing task that can be easily adapted to documentation projects, for both seen languages and unseen ones.
 6.1 Architecture
-We base GlossLM off of ByT5, a multilingual pretrained model using the T5 architecture Raffel et al.
+We use the ByT5 model, a multilingual pretrained model using the T5 architecture Raffel et al.
 (2020).
@@
 6.2 IGT Pretraining
-We continually pretrain ByT5 on two versions of the train split of our dataset.
-The first (GlossLMall) uses all the training data in our dataset.
-Our second model (GlossLMunseg) excludes any segmented data in any of our test languages, simulating scenarios where a linguist does not have access to a corpus of segmented glossed data in their target language.
-Then, for each version of GlossLM, we finetune on the target language of interest, again using all data in the target language for GlossLMall and only unsegmented data in the target language for GlossLMunseg.
-As ByT5 is based on the T5 model, we cast the glossing task into a text-to-text format. We train the model using examples formatted with the following prompt:
+We continually pretrain the ByT5 model on the GlossLM corpus described in section 3.
+As we are evaluating on unsegmented IGT data, we omit segmented data for the evaluation languages from the pretraining corpus.
+[^3] However, our corpus includes segmented IGT examples in other languages, which we do not evaluate.
+We structure the glossing task as a text-to-text problem, training the model with examples formatted with the following prompt:
 Provide the glosses for the following transcription in <lang>.
@@
 Models are trained to output the gloss line following the above prompt input.
-Including translations has been shown to help in predicting glosses (Zhao et al., 2020; Ginn et al., 2023).
+We include translations, which has been shown to provide benefits in gloss prediction (Zhao et al., 2020; Ginn et al., 2023).
 For some data, a translation was not available ($\approx 3\%$ of the training data), in which case the translation line is omitted.
@@
 We did not conduct hyperparameter search, only tuning the batch size, to fit in our GPUs, and epochs and early stopping, to ensure convergence.
-5 Experiments
-We evaluate the effectiveness of our pretrained model on automatic generation of IGT across the seven evaluation languages described in §3.4.
-5.1 Experimental Settings
-We compare models under two settings.
-Segmented
-In the segmented setting, words in the input transcriptions are divided into morphemes with segmentation boundaries.
-Models must learn to classify each morpheme with an appropriate gloss, making this an easier setting (with SOTA models achieving 80-90% accuracy).
-Unsegmented
-In the unsegmented setting, words in the input transcriptions are not segmented (but glosses are), requiring the model to jointly learn segmentation and glossing.
-Generally, this setting is far more difficult, with SOTA models achieving as low as 11% accuracy on small datasets.
-5.2 GlossLM Experiments
-We evaluate the GlossLM model in two variations.
-First, we directly evaluate the pretrained model on the in-domain languages without any additional finetuning, in order to analyze whether the model can retain high performance across individual languages despite the highly multilingual pretraining corpus.
-We refer to these models as GlossLMpre.
-We then finetune the GlossLM pretrained model on the training dataset for each language individually, and label these runs as GlossLMft.
-We do so for both the in-domain languages, to focus the model on the target language, as well as the out-of-domain languages, allowing us to study the model’s adaptation to unseen languages.
-Finetuning used the same parameters as pretraining, but with 100 training epochs and early stopping (patience 3, start epoch 15) [^3] For Gitksan, due to the size of the training set, we set max epochs to 300 and patience to 15., and took anywhere from 20 minutes to around 4 hours for each language.
+6.3 Performance of Pretrained Model
+When training massively multilingual models, performance on individual languages can sometimes degrade in what is dubbed the "curse of multilinguality" (Conneau et al., 2020; Chang et al., 2023).
+To investigate this issue, we evaluate our pretrained model on the in-domain languages without any additional finetuning.
+We compare the performance of our pretrained model to the current SOTA, which is the second system from Girrbach (2023a), as shown in Figure 3.
+We find that the model outperforms the SOTA across all three in-domain languages.
+This result gives little evidence to believe our model suffers from the curse of multilinguality, as it retains good performance across several languages.
+[^4] These languages do make up large fractions of our pretraining corpus, so the model will almost certainly underperform on underrepresented languages.
+Figure 3: Comparison of our pretrained model and the SOTA Girrbach (2023a) for in-domain languages on unsegmented data. Our model outperforms on all three languages.
+Our pretrained model can be used for automated glossing across several languages, without needing to train and serve separate monolingual models.
+This could be valuable to real-world documentation projects, as we can serve a single pretrained model that can be used across projects, significantly reducing the barrier to using an automated system.
+7 Results
+7.1 Comparison with Baselines
+Figure 4: Morpheme accuracy for various systems.
+After pretraining GlossLM, we train finetuned versions for each of the languages in the test set.
+We first describe our finetuning procedure and compare results of our finetuned models against baselines from Ginn (2023)7.1).
+Then, to further isolate the efficacy of pretraining, we compare the finetuned versions of GlossLM to a finetuned ByT5 model without the multilingual gloss pretraining (§7.2).
+Finally, we explore whether training on a minimally normalized version of the data improves performance (§7.3).
+As previously noted, we focus on the unsegmented setting; for completeness, we provide full results for both segmented and unsegmented data in Appendix D.
+Finetuning can help align the model to a particular language or even a new unseen language.
+We finetune our GlossLM pretrained model on the training dataset for each language individually, and label these runs as GlossLMft.
+We do this for both the in-domain languages, to focus the model on a single language, as well as the out-of-domain languages, allowing us to study the model’s adaptation to unseen languages.
+Finetuning used the same parameters as pretraining, but with 100 training epochs and early stopping (patience 3, start epoch 15) [^5] For Gitksan, due to the size of the training set, we set max epochs to 300 and patience to 15., and took anywhere from 20 minutes to one day for each language.
 Inference uses beam search with $n=3$ beams.
-5.3 Baseline Experiments
-We compare the GlossLM models with six baseline systems which include the state-of-the-art from prior work.
-Top-choice
-selects the most frequent label associated with each morpheme/word in the training data, and assigns “???” to unseen morphemes.
-Token-class
-treats the glossing task as a token classification problem, where the output vocabulary consists of the IGT morpheme or word-level labels.
+We compare the finetuned GlossLM models with three baseline systems which include the state-of-the-art from prior work:
+  1. 1.
+    Top-choice selects the most frequent label associated with each morpheme/word in the training data, and assigns “???” to unseen morphemes.
+  2. 2.
+    Token-class treats the glossing task as a token classification problem, where the output vocabulary consists of the IGT morpheme or word-level labels.
 Each target language’s data is used to train a language-specifc Token-class model, which uses the RoBERTa architecture with default hyperparameters without any additional pretraining Liu et al.
@@
 This was used as the baseline model for the SIGMORPHON 2023 Shared Task on interlinear glossing Ginn (2023).
-Tü-CL Girrbach (2023b)
-uses straight-through gradient estimation (Bengio et al., 2013) to induce latent, discrete segmentations of input texts, and predicted glosses using a multilayer perceptron.
-CRF Okabe and Yvon (2023b)
-is a hybrid CRF-neural system that is composed of an unsupervised neural alignment system (SimAlign, Jalili Sabet et al.
-2020) to make use of lexemes in translations and a CRF sequence labeling system (Lost, Lavergne et al.
-2011).
-SigMoreFun (SMF) He et al. (2023)
-also continually pretrain ByT5, though on a smaller amount of data (using the ODIN corpus). Data from all target languages were combined and shuffled for finetuning.
-ByT5all and ByT5unseg
-are ByT5 models Xue et al.
-(2022) finetuned on the target language data without any multilingual continual pretraining on IGT.
-We otherwise replicate the finetuning procedure described in subsection 5.2.
-5.4 Evaluation
-For evaluating predictions, we strip punctuation (except for within glosses).
-[^4] Because of this post-processing, our results for baselines are slightly different than what original sources report.
-We evaluate morpheme accuracy, which counts the number of correct morpheme glosses in the correct position.
-Hence, if a gloss is inserted or deleted, the following glosses will be incorrect.
-We also evaluate word accuracy, which counts the number of entire correct word glosses.
-However, accuracy may sometimes be too strict of a measurement—especially for generative models—as minor character insertions/deletions in the label are penalized heavily.
-Thus, we also evaluate chrF++ (Popović, 2015), a character-level metric often used in machine translation.
-chrF++ measures the F1 score over character n-grams between the reference and predictions, and is robust to insertions and deletions, unlike accuracy.
-Model Morpheme accuracy / Word accuracy (Segmented)
-In-domain Out-of-domain
-arp ddo usp Avg git lez ntu nyb Avg
-Top-choice 83.2 / 74.0 78.5 / 64.4 79.7 / 72.9 80.5 / 70.4 51.1 / 29.7 62.2 / 54.4 78.4 / 68.1 72.5 / 63.8 66.1 / 54.0
-Token-class 90.4 / 84.2 86.4 / 76.5 82.5 / 76.5 86.4 / 79.1 25.3 / 16.4 50.2 / 38.8 62.0 / 54.3 88.8 / 84.4 56.6 / 48.5
-Tü-CL 90.7 / 84.6 91.2 / 85.1 84.5 / 78.5 88.8 / 82.7 50.2 / 26.6 84.9 / 77.6 91.7 / 86.0 91.4 / 87.9 79.6 / 69.5
-CRF 90.4 / 84.2 91.9 / 85.6 84.4 / 78.9 88.9 / 82.9 52.4 / 33.6 84.7 / 77.5 91.1 / 86.6 88.8 / 84.4 79.3 / 70.5
-SMF 80.1 / 79.4 78.2 / 82.8 73.2 / 75.7 77.2 / 79.3 12.7 / 20.6 47.8 / 56.4 64.0 / 75.7 85.4 / 82.7 52.5 / 58.9
-ByT5all 88.7 / 83.2 93.5 / 89.9 86.3 / 82.7 89.5 / 85.3 2.2 / 3.6 72.5 / 69.7 83.4 / 82.2 90.7 / 89.2 62.2 / 61.2
-GlossLMall, pre 89.3 / 84.2 91.7 / 88.3 84.1 / 81.0 88.4 / 84.5 3.6 / 9.1 3.6 / 1.8 4.9 / 9.8 2.9 / 3.0 3.8 / 5.9
-GlossLMall, ft 90.1 / 85.0 92.8 / 89.3 86.4 / 84.5 89.8 / 86.3 28.9 / 34.9 74.7 / 71.3 86.0 / 81.5 90.7 / 87.7 70.1 / 68.9
-Model Morpheme accuracy / Word accuracy (Unsegmented)
-In-domain Out-of-domain
-arp ddo usp Avg git lez ntu nyb Avg
-Top-choice 27.9 / 56.9 15.2 / 64.1 43.6 / 60.4 28.9 / 60.5 3.6 / 16.9 20.1 / 58.2 12.7 / 55.1 72.3 / 76.7 27.2 / 51.7
-Token-class 43.6 / 69.9 51.2 / 74.3 57.2 / 72.1 50.7 / 72.1 8.54 / 16.9 40.7 / 45.5 19.4 / 48.2 14.2 / 5.96 20.7 / 29.1
-Tü-CL 77.8 / 77.5 74.1 / 80.4 70.0 / 73.4 74.0 / 77.1 11.7 / 21.1 59.9 / 71.8 56.2 / 78.0 85.2 / 85.0 53.3 / 64.0
-ByT5unseg 80.8 / 79.7 84.2 / 87.4 78.9 / 82.5 81.3 / 83.2 0.1 / 0.3 42.2 / 53.4 53.7 / 71.0 90.4 / 88.4 46.6 / 53.3
-GlossLMunseg, pre 79.8 / 79.2 77.5 / 82.8 76.8 / 80.8 78.0 / 80.9 2.3 / 3.5 1.5 / 1.3 4.1 / 9.6 1.6 / 2.9 2.4 / 4.3
-GlossLMunseg, ft 82.1 / 81.5 83.6 / 87.3 78.6 / 81.0 81.4 / 83.3 10.1 / 28.4 57.3 / 64.9 62.8 / 78.9 87.4 / 86.2 56.3 / 65.0
-Table 3: Morpheme- and word-level accuracy of various systems on segmented (top) and unsegmented (bottom) text.
-Best performance per language in each setting the table is bolded.
-GlossLMall, pre refers to performance using the pretrained GlossLM directly, while GlossLMall, ft refers to performance after fine-tuning the pretrained model on the specific language.
-6 Results
-Figure 2: Word accuracy for the multilingual GlossLM and various baseline systems on unsegmented data for out-of-domain languages.
-Accuracy results for the segmented setting are given in Table 3, and for the unsegmented setting in Table 3 and Figure 2.
-We report chrF++ scores in Appendix D, finding that they generally show the same trends as accuracy.
-[^5] All results are for a single run, due to the computational cost of multiple runs.
-Segmented setting
-In the segmented setting, we observe mixed results.
-The GlossLM models outperform baselines for certain languages (arp, usp, git) but underperform in other cases (ddo, lez, ntu, nyb).
-GlossLM matches or outperforms the comparable ByT5 baseline, indicating that the continual gloss pretraining provides small benefits.
-However, this method does not necessarily overcome other architectures that are more successful, such as the straight-through gradient model Girrbach (2023a) (though future work could utilize our corpus to pretrain models with these architectures).
-We note that for this setting, performance across models is very high, and it may be difficult to achieve much higher than the observed performances (due to ambiguous morphemes or noisy evaluation data).
-Notably, the direct ByT5 model trained on Gitksan far underfits the data, producing improperly formatted data, long random sequences, and outputs including chunks of the prompts.
-As the Gitksan dataset is very small, the model was unable to learn the correct formatting; meanwhile, the GlossLM model benefits from learning correct formatting from other languages, for a 26.7% morpheme accuracy improvement.
-For very low-resource documentation projects, our pretrained model shows clear benefits in producing outputs with the correct formatting, while leveraging the multilingual knowledge of ByT5.
-We also observe that while finetuning on in-domain languages provides small improvements in performance, the non-finetuned models perform very well.
-This indicates that multilinguality has not limited performance on each individual language too much, despite the hugely multilingual pretraining corpus.
-Furthermore, the non-finetuned model is able to make a small percentage of correct predictions (2.9% to 4.9% morpheme accuracy) on unseen languages without training, indicating clear benefits from massively multilingual pretraining.
-Unsegmented setting
-In the unsegmented setting, the GlossLM models outperform most prior baselines on the in-domain data, although they slightly underperform the ByT5 baseline.
-This result indicates that with large monolingual datasets, there is not necessarily a benefit to additional multilingual learning on IGT data.
-This could indicate that the multilingual (non-IGT) corpus used to pretrain ByT5 already sufficiently provides any benefits from crosslingual transfer.
-The non-finetuned model is also highly competitive, beating prior baselines (except the ByT5 baseline).
-For the out-of-domain languages (which tend to have much smaller corpora), the finetuned GlossLM model outperforms the ByT5 baseline on three of four languages and beats previous best baselines by up to 6.6% morpheme accuracy (for Natugu).
-When the monolingual dataset is very limited, there seems to be a greater benefit from multilingual IGT pretraining, allowing the model to leverage information about segmentation and output formatting from languages in the pretraining corpus.
-The only case where the finetuned model underperforms SOTA is for Lezgi, slightly underperforming the Tü-CL model.
-One possibility is that segmentation from surface forms in Lezgi is more difficult, and the hard attention mechanism is beneficial at learning proper morpheme boundaries.
-Overall, we suggest the pretrained GlossLM model is a good fit for languages with limited training data, even unseen languages.
-Directly finetuning a ByT5 model can be very effective for languages with large training corpora, but few documentation projects will have as much available data as the in-domain languages used here (which have 7k to 78k training sentences), and finetuning directly on small datasets risks failure to converge.
-Furthermore, we observe that finetuning the GlossLM model on unseen data converges much faster than ByT5–for example, after 50 train epochs for Natugu, the GlossLM-based model achieves 78.1 chrF++ on the eval set, while the ByT5 model is at 30.1.
-If the desire is to conserve computational resources, using the pretrained model could be beneficial to train efficiently.
-7 Error Analysis
-7.1 In- vs out-of-vocabulary errors
-In-domain Out-of-domain
-arp ddo usp git lez ntu nyb
-% OOV 30.0 15.6 20.0 78.1 27.3 27.6 8.42
-IV 96.2 92.3 91.4 66.7 85.4 91.3 92.8
-OOV 55.7 71.7 57.1 26.0 33.9 55.7 32.6
-% OOV 30.0 18.7 21.4 80.5 25.5 28.9 9.27
-IV 95.3 91.4 89.1 60.0 81.6 91.3 91.7
-OOV 50.1 69.7 50.9 20.7 16.4 48.3 30.6
-Table 4: Percent of words that are out-of-vocab in the test split for each language along with in- versus out-of-vocabulary accuracy at the word level.
-Top is the segmented setting (GlossLMall, ft), bottom is unsegmented (GlossLMunseg, ft).
-We are interested in how well our models can generalize to unseen words, which are often novel combinations of new and previously-seen morphemes.
-We report word-level accuracy for our finetuned GlossLM models, indicating whether the transcribed word is in- or out-of-vocabulary in Table 4, as well as the percent of OOV words in the test set.
-The OOV rate between segmented and unsegmented may vary slightly, as mappings between segmented and unsegmented forms are not necessarily one-to-one.
-We consider a word to be in-vocabulary if the form of the word in the transcription and its corresponding gold label in the gloss co-occur in the training data.
-We also include morpheme-OOV rates and statistics on lexical gloss overlap with translations in Appendix E.
-Across most languages, our models perform well on in-vocabulary words (with Gitksan being a notable exception, likely due to the small amount of training data).
-Within the OOV split, we expect that models perform better on languages that are agglutinative with lower morpheme-OOV rates.
-We find that this is the case, as Tsez (which has the highest OOV accuracy) is a highly agglutinative language with the lowest rate of OOV morphemes in the test set, implying that the majority of OOV words are actually composed of seen morphemes.
-However, Lezgi—which is also highly agglutinative—has strikingly low OOV accuracy; this may be due to the highly complex inflections for case marking, which could be more difficult for our system to learn.
-Finally, Nyangbo also shows relatively low OOV word accuracy, possibly due to the lack of translations in the input, which could help predict OOV stem labels.
-7.2 Qualitative Analysis of Errors
-A manual analysis of GlossLM outputs reveal two interesting points of consideration.
-In certain cases, the model may rely too heavily on translations when predicting lexical glosses.
-When inspecting the Uspanteko outputs from the finetuned GlossLMall model, we found that the model outputs “esta@bien” (the correct label for the morpheme tziin) as the label for jaan, which should be glossed as “bueno”; this happens especially often when the phrase “esta bien” appears in the translation.
-In other cases, the model sometimes outputs a word or phrase present in the translation, even if there is no occurrence of that item in the gold gloss.
-We include example system outputs to illustrate these cases in Appendix F.
-Our metrics potentially underestimate GlossLM’s performance on the glossing task.
-There are sometimes inconsistencies in the IGT labels where multiple interchangeable glosses are used for the same morpheme, particularly for lexical glosses (e.g.
-the Arapaho word ’eeneisih’i is glossed in the data as “how.X.things.are.named”, “how.s.o..is.named”, and “how.things.are.named”).
-These types of errors can be accounted for when using chrF++, but another issue that may more greatly impact generative models—especially if it learns to use translations effectively—is if the model outputs a semantically similar label that is dissimilar in form; for example, outputting “PREP” (preposition label) instead of “en” (“in” in Spanish, which is a preposition).
-We include examples of such outputs in Appendix F.
+  3. 3.
+    Tü-CL Girrbach (2023b) uses straight-through gradient estimation (Bengio et al., 2013) to induce latent, discrete segmentations of input texts, and predicts glosses using a multilayer perceptron.
+As shown in  Figure 4, we find that our finetuned models outperform SOTA in all but two languages (Gitksan and Lezgi).
+The Tü-CL model (Girrbach, 2023a), which is a close second and outperforms on Gitksan and Lezgi, uses explicit latent segmentations, which seems to be particularly beneficial for the very low-resource, unsegmented setting.
+7.2 Comparison with Finetuned ByT5
+To directly assess the impact of pretraining on performance, we finetune ByT5 models on each language in the test set with the same configuration as for the finetuned GlossLM models.
+We then compare the performance of our models (which have under gone both multilingual gloss pretraining and finetuning) with analogous ByT5 models (without multilingual gloss pretraining), as shown in Figure 5.
+Figure 5: Performance after monolingual finetuning, comparing a standard pretrained ByT5 with a continually pretrained GlossLM model.
+The x-axis uses the log (base 10) of the number of training examples in a given language, for readability.
+We observe mixed results, which are largely dependent on the size of the training corpus.
+For languages with less training data (Gitksan, Lezgi, Natugu) the GlossLMft model outperforms the finetuned ByT5 model (by 10.0, 15.1, and 9.1 points respectively).
+In the case of Gitksan, the finetuned ByT5 model is completely unable to produce well-formatted output (likely due to the tiny training corpus) while the GlossLMft model does not struggle with this as much.
+A possible explanation is that even if there are no similar languages in the pretraining corpus, the GlossLMft can leverage knowledge about IGT formatting from unrelated languages.
+With enough data (starting with Nyangbo, 2100 examples), the two approaches achieve nearly identical performance.
+This is an unsurprising result, indicating that large amounts of high-quality monolingual data overshadow any benefits from crosslingual transfer.
+Furthermore, we note that benefits of multilingual gloss pretraining shown may not be unique to the T5 architecture—while we only experiment with ByT5, our pretraining strategy could be applied to other architectures.
+7.3 Effect of Gloss Normalization
+Finally, we experiment with pretraining and finetuning on a minimally normalized version of the dataset, where the 200 most frequent grammatical labels are mapped to a set of standard labels.
+We repeat the same pretraining and finetuning process as before.
+When comparing the performance of the pretrained model before finetuning on in-domain languages, we find minimal differences compared to the results in §6.3.
+This is also the case after finetuning, though we found improvements in three of the four out-of-domain languages: Lezgi ($+\Delta 2.9$), Natugu ($+\Delta 6.1$), and Nyangbo ($+\Delta 2.5$), which (aside from Gitksan) are the three lowest resourced languages in our test set, exceeding current SOTA performance.
+From this, it appears that normalization may be most helpful when adapting the pretrained model to an unseen language with a low-to-moderate amount of training data.
 8 Related Works
-8.1 Automatic Interlinear Glossing
+Automatic Interlinear Glossing
 Recent research has explored various methods for generating IGT, including rule-based methods (Bender et al., 2014), active learning (Palmer et al., 2010, 2009), conditional random fields (Moeller and Hulden, 2018; McMillan-Major, 2020), and neural models (Moeller and Hulden, 2018; Zhao et al., 2020).
@@
 Furthermore, neither of these studies find significant benefits to using pretraining corpora, while we observe benefits under certain conditions.
-8.2 Large Multilingual Pretrained Models
+Large Multilingual Pretrained Models
 Prior work has shown that large, massively multilingual pretrained language models can boost performance across low- and high-resource languages on a variety of tasks.
@@
 9 Conclusion
-Automatic generation of interlinear glossed text has great potential to aid documentation projects, but limited monolingual data limits effectiveness.
-In this research, we attempt to overcome these limitations using crosslingual transfer, compiling the largest multilingual IGT dataset and pretraining massively multilingual glossing models.
-Our models outperform any existing work in the majority of settings, demonstrating clear benefits of multilingual representations–particularly for languages with very small corpora.
-We make our dataset and models available for future research, in hopes that they will aid future NLP research and benefit documentation projects.
-10 Limitations
+High-quality language documentation involves an incredible amount of effort.
+We compile, normalize, and release the largest corpus of multilingual IGT data, enabling future research in linguistics, NLP, and documentation.
+Furthermore, we demonstrate the applicability of our corpus by pretraining a multilingual neural model for automatic generation of IGT.
+We finetune the model on monolingual corpora, showing benefits on low-resource languages due to multilingual pretraining.
+In five out of seven languages, we achieve a new SOTA on automatic IGT generation.
+Limitations
 Our work evaluates the effectiveness of massively multilingual pretraining on seven IGT datasets in different languages using the ByT5 architecture.
@@
 Our models will likely have impacted performance for out-of-domain texts, such as highly technical or domain-specific language.
-11 Ethical Considerations
+Ethics Statement
 We hope this work can aid in the struggle against language extinction.
@@
 Table 4: Pretraining Hyperparameters
-Appendix D chrF++ Scores
-We provide chrF++ scores for various systems in Table 10. Broadly, we found the results supported the same trends as the accuracy measurements, although the magnitude of the differences varied.
-Appendix E Test Split Vocabulary Statistics
+Appendix D Full Results
+We provide full results for accuracy and chrF++ scores in Table 5 and Table 6 . For our normalization experiments, we only trained and tested on unsegmented data for the target languages.
+Model Morpheme accuracy / Word accuracy (Segmented)
+In-domain Out-of-domain
+arp ddo usp Avg git lez ntu nyb Avg
+Top-choice 83.2 / 74.0 78.5 / 64.4 79.7 / 72.9 80.5 / 70.4 51.1 / 29.7 62.2 / 54.4 78.4 / 68.1 72.5 / 63.8 66.1 / 54.0
+Token-class 90.4 / 84.2 86.4 / 76.5 82.5 / 76.5 86.4 / 79.1 25.3 / 16.4 50.2 / 38.8 62.0 / 54.3 88.8 / 84.4 56.6 / 48.5
+Tü-CL 90.7 / 84.6 91.2 / 85.1 84.5 / 78.5 88.8 / 82.7 50.2 / 26.6 84.9 / 77.6 91.7 / 86.0 91.4 / 87.9 79.6 / 69.5
+CRF 90.4 / 84.2 91.9 / 85.6 84.4 / 78.9 88.9 / 82.9 52.4 / 33.6 84.7 / 77.5 91.1 / 86.6 88.8 / 84.4 79.3 / 70.5
+SMF 80.1 / 79.4 78.2 / 82.8 73.2 / 75.7 77.2 / 79.3 12.7 / 20.6 47.8 / 56.4 64.0 / 75.7 85.4 / 82.7 52.5 / 58.9
+ByT5all 88.7 / 83.2 93.5 / 89.9 86.3 / 82.7 89.5 / 85.3 2.2 / 3.6 72.5 / 69.7 83.4 / 82.2 90.7 / 89.2 62.2 / 61.2
+GlossLMall, pre 89.3 / 84.2 91.7 / 88.3 84.1 / 81.0 88.4 / 84.5 3.6 / 9.1 3.6 / 1.8 4.9 / 9.8 2.9 / 3.0 3.8 / 5.9
+GlossLMall, ft 90.1 / 85.0 92.8 / 89.3 86.4 / 84.5 89.8 / 86.3 28.9 / 34.9 74.7 / 71.3 86.0 / 81.5 90.7 / 87.7 70.1 / 68.9
+Model Morpheme accuracy / Word accuracy (Unsegmented)
+In-domain Out-of-domain
+arp ddo usp Avg git lez ntu nyb Avg
+Top-choice 27.9 / 56.9 15.2 / 64.1 43.6 / 60.4 28.9 / 60.5 3.6 / 16.9 20.1 / 58.2 12.7 / 55.1 72.3 / 76.7 27.2 / 51.7
+Token-class 43.6 / 69.9 51.2 / 74.3 57.2 / 72.1 50.7 / 72.1 8.54 / 16.9 40.7 / 45.5 19.4 / 48.2 14.2 / 5.96 20.7 / 29.1
+Tü-CL 77.8 / 77.5 74.1 / 80.4 70.0 / 73.4 74.0 / 77.1 11.7 / 21.1 59.9 / 71.8 56.2 / 78.0 85.2 / 85.0 53.3 / 64.0
+ByT5unseg 80.8 / 79.7 84.2 / 87.4 78.9 / 82.5 81.3 / 83.2 0.1 / 0.3 42.2 / 53.4 53.7 / 71.0 90.4 / 88.4 46.6 / 53.3
+GlossLMunseg, pre 79.8 / 79.2 77.5 / 82.8 76.8 / 80.8 78.0 / 80.9 2.3 / 3.5 1.5 / 1.3 4.1 / 9.6 1.6 / 2.9 2.4 / 4.3
+GlossLMunseg, ft 82.1 / 81.5 83.6 / 87.3 78.6 / 81.0 81.4 / 83.3 10.1 / 28.4 57.3 / 64.9 62.8 / 78.9 87.4 / 86.2 54.4 / 64.6
+GlossLM-normunseg, pre 79.6 / 80.0 79.6 / 83.2 74.8 / 76.6 78.0 / 79.9 2.2 / 7.8 2.6 / 1.8 2.9 / 9.7 1.0 / 2.5 2.2 / 5.45
+GlossLM-normunseg, ft 82.0 / 81.5 84.2 / 87.8 76.4 / 79.2 80.8 / 82.8 9.3 / 16.4 60.3 / 67.8 63.4 / 76.6 90.0 / 87.6 55.8 / 62.1
+Table 5: Morpheme- and word-level accuracy of various systems on segmented (top) and unsegmented (bottom) text.
+Best performance per language in each setting the table is bolded.
+GlossLMall, pre refers to performance using the pretrained GlossLM directly, while GlossLMall, ft refers to performance after fine-tuning the pretrained model on the specific language.
+Model chrF++ Score (Segmented)
+In-domain Out-of-domain
+arp ddo usp Avg git lez ntu nyb Avg
+Top-choice 75.0 71.9 71.4 72.8 33.7 69.7 74.5 62.2 60.0
+Token-class 84.2 81.8 75.3 80.4 25.6 52.1 65.4 84.3 56.9
+Tü-CL 85.2 88.4 77.7 83.8 34.8 78.9 87.9 87.7 72.3
+CRF 84.2 88.2 79.4 83.9 40.9 79.2 88.4 84.3 73.2
+SMF 80.7 86.6 76.3 81.2 28.0 62.0 78.8 82.2 62.6
+ByT5all 84.2 91.8 84.6 86.9 9.07 74.9 85.1 88.8 64.5
+GlossLMall, pre 85.2 90.8 83.1 86.4 21.5 14.1 18.2 8.8 15.7
+GlossLMall, ft 86.3 91.5 85.9 87.9 43.1 75.0 84.1 87.4 72.4
+Model chrF++ Score (Unsegmented)
+In-domain Out-of-domain
+arp ddo usp Avg git lez ntu nyb Avg
+Top-choice 44.0 63.5 55.1 54.2 8.4 51.6 40.5 74.0 43.6
+Token-class 56.2 72.9 65.3 64.8 18.8 56.4 45.1 18.8 34.8
+Tü-CL 77.6 84.6 72.5 78.2 23.0 71.5 78.6 84.1 64.3
+ByT5unseg 80.7 90.0 83.0 84.6 7.6 59.6 77.0 88.4 58.2
+GlossLMunseg, pre 80.5 86.8 81.0 82.8 19.4 13.3 16.3 8.1 14.3
+GlossLMunseg, ft 82.9 89.8 81.7 84.8 34.9 68.8 80.7 85.5 67.5
+GlossLM-normunseg, pre 80.3 86.6 78.7 81.9 19.5 14.3 16.6 7.4 14.5
+GlossLM-normunseg, ft 82.6 89.9 81.1 84.5 29.1 70.9 79.0 87.5 66.6
+Table 6: chrF++ scores of various systems on segmented (top) and unsegmented (bottom) data.
+Best performance per language in each setting the table is bolded.
+GlossLMall, pre refers to performance using the pretrained GlossLM directly, while GlossLMall, ft refers to performance after finetuning on the specific language.
+Appendix E In- vs out-of-vocabulary errors
+In-domain Out-of-domain
+arp ddo usp git lez ntu nyb
+% OOV 30.0 15.6 20.0 78.1 27.3 27.6 8.42
+IV 96.2 92.3 91.4 66.7 85.4 91.3 92.8
+OOV 55.7 71.7 57.1 26.0 33.9 55.7 32.6
+% OOV 30.0 18.7 21.4 80.5 25.5 28.9 9.27
+IV 95.3 91.4 89.1 60.0 81.6 91.3 91.7
+OOV 50.1 69.7 50.9 20.7 16.4 48.3 30.6
+Table 7: Percent of words that are out-of-vocab in the test split for each language along with in- versus out-of-vocabulary accuracy at the word level.
+Top is the segmented setting (GlossLMall, ft), bottom is unsegmented (GlossLMunseg, ft).
+We report word-level accuracy for our finetuned GlossLM models, indicating whether the transcribed word is in- or out-of-vocabulary in Table 7, as well as the percent of OOV words in the test set.
+The OOV rate between segmented and unsegmented may vary slightly, as mappings between segmented and unsegmented forms are not necessarily one-to-one.
+We consider a word to be in-vocabulary if the form of the word in the transcription and its corresponding gold label in the gloss co-occur in the training data.
+We also include morpheme-OOV rates and statistics on lexical gloss overlap with translations in in Table 8 and Table 9, as reported in Ginn et al.
+(2023).
 Language Morpheme %OOV
@@
 Table 9: Percent of lexical glosses present in the translation in the test split. Nyangbo examples do not include translations.
-We include additional statistics for morpheme OOV rate and recall rate for lexical labels in the translation in Table 7 and Table 8, as reported in Ginn et al. (2023).
 Appendix F Example System Outputs
@@
 Table 10: Selected example outputs to illustrate errors by GlossLMall finetuned models.
-We include examples of system outputs to illustrate the types of errors discussion in subsection 7.2 in Table 9.
-Model chrF++ Score (Segmented)
-In-domain Out-of-domain
-arp ddo usp Avg git lez ntu nyb Avg
-Top-choice 75.0 71.9 71.4 72.8 33.7 69.7 74.5 62.2 60.0
-Token-class 84.2 81.8 75.3 80.4 25.6 52.1 65.4 84.3 56.9
-Tü-CL 85.2 88.4 77.7 83.8 34.8 78.9 87.9 87.7 72.3
-CRF 84.2 88.2 79.4 83.9 40.9 79.2 88.4 84.3 73.2
-SMF 80.7 86.6 76.3 81.2 28.0 62.0 78.8 82.2 62.6
-ByT5all 84.2 91.8 84.6 86.9 9.07 74.9 85.1 88.8 64.5
-GlossLMall, pre 85.2 90.8 83.1 86.4 21.5 14.1 18.2 8.8 15.7
-GlossLMall, ft 86.3 91.5 85.9 87.9 43.1 75.0 84.1 87.4 72.4
-Model chrF++ Score (Unsegmented)
-In-domain Out-of-domain
-arp ddo usp Avg git lez ntu nyb Avg
-Top-choice 44.0 63.5 55.1 54.2 8.4 51.6 40.5 74.0 43.6
-Token-class 56.2 72.9 65.3 64.8 18.8 56.4 45.1 18.8 34.8
-Tü-CL 77.6 84.6 72.5 78.2 23.0 71.5 78.6 84.1 64.3
-ByT5unseg 80.7 90.0 83.0 84.6 7.6 59.6 77.0 88.4 58.2
-GlossLMunseg, pre 80.5 86.8 81.0 82.8 19.4 13.3 16.3 8.1 14.3
-GlossLMunseg, ft 82.9 89.8 81.7 84.8 34.9 68.8 80.7 85.5 67.5
-Table 10: chrF++ scores of various systems on segmented (top) and unsegmented (bottom) data.
-Best performance per language in each setting the table is bolded.
-GlossLMall, pre refers to performance using the pretrained GlossLM directly, while GlossLMall, ft refers to performance after finetuning on the specific language.
+We include examples of system outputs to illustrate the types of errors discussion in Table 10.
 ID Avg. Value ID (cont.) Avg. Value (cont.) ID (cont.) Avg. Value (cont.)

### v3

@@
 We demonstrate the utility of this model by finetuning it on monolingual corpora, outperforming SOTA models by up to 6.6%.
-We will make our pretrained model and dataset available through Hugging Face, as well as provide access through a web interface for use in language documentation efforts.
+Our pretrained model and dataset are available on Hugging Face.
+[^1] https://huggingface.co/collections/lecslab/glosslm-66da150854209e910113dd87
 […]
@@
 To address this, we compile the largest corpus of digitized IGT from various existing sources, with over 450k examples in 1.8k languages (§3).
-We explore methods to normalize this data (§4)., standardizing over 80% of the grammatical glosses in the corpus to follow the UniMorph (Sylak-Glassman, 2016).
-We will release our corpora for future NLP, linguistics, and language documentation research.
+We explore methods to normalize this data (§4)., standardizing over 80% of the grammatical glosses in the corpus to follow the UniMorph schema (Sylak-Glassman, 2016).
+We are releasing our corpora for future NLP, linguistics, and language documentation research.
 Automating IGT Generation
-Futhermore, creating new IGT corpora is difficult and time-consuming, requiring documenters to perform linguistic analysis and extensive documentation simultaneously.
+The creation of new IGT corpora is often difficult and time-consuming, requiring documenters to perform linguistic analysis and extensive documentation simultaneously.
 Research has found that computational tools can help accelerate annotation and overcome this bottleneck Palmer et al.
@@
 We aim to overcome the monolingual data bottleneck by creating a multilingual pretrained glossing model that can be adapted to specific languages and gloss formats.
-We continually pretrain a model on our corpus, finding that pretrained multilingual model retains high performance across languages.
+We continually pretrain a model on our corpus, and find that the pretrained multilingual model retains high performance across languages.
 We then finetune the pretrained model on monolingual data, including languages that do not appear in the pretraining corpus.
-We achieve a new SOTA performance on five out of seven languages, and demonstrate clear improvements for low-resource language settings over an equivalent finetuned model without our continual pretraining (§7).
+Our models achieve new SOTA performance on five out of seven languages, and demonstrate clear improvements for low-resource language settings over an equivalent finetuned model without our continual pretraining (§7).
 2 Interlinear Glossed Text (IGT)
@@
 3 GlossLM Corpus
-We compile and clean the largest IGT dataset from a variety of sources and languages.
+While various publicly available sources of digitized IGT exist, the lack of unified data formatting and ease of access is a roadblock to using these resources effectively.
+To solve this problem, we compile and clean the largest IGT dataset from a variety of sources and languages.
 In total, our dataset contains over 450k IGT instances (from 250k unique sentences) in 1.8k languages, collected from six different IGT corpora.
@@
 4.1 Motivation
-As our data comes from a variety of sources, spanning many languages and documentation projects, there is a great amount of diversity of the morphological glosses used.
+As our data comes from a variety of sources, spanning many languages and documentation projects, there is a great amount of diversity in the morphological glosses used.
 This includes cases where several different labels are used to indicate the same feature (e.g.
@@
 1Sg vs 1.Sg).
-We explore the value and feasibility of normalizing glosses to a single standardized format.
-On one hand, normalizing glosses may make it easier to train models that utilize crosslingual information through shared gloss labels, but on the other hand, it can be difficult to select a single schema that preserves the original intent of all annotators.
-We split gloss lines by period and count the number and frequency of unique grammatical gloss labels across our corpus (focusing on the all-caps functional glosses, not stems) and visualize the distribution in Figure 2.
+We explore the feasibility and value of normalizing glosses to a single standardized format.
+On one hand, normalizing glosses may make it easier to train models that utilize crosslingual information through shared gloss labels, but on the other hand, it is difficult (perhaps impossible) to select a single schema that preserves the original intent of all annotators.
+We split gloss lines by period and count the number and frequency of unique grammatical gloss labels across our corpus (focusing on the all-caps functional glosses, not stem translations) and visualize the distribution in Figure 2.
 Figure 2: Distribution of unique glosses across all languages.
@@
 For example, the label S (appearing 20,855 times) is used for singular, subject, and noun/sustantivo (at least) in our dataset.
-In order to map these, we would need to analyze on a sentence-by-sentence basis, which was not feasible; thus, we left these labels as-is.
+In order to map these, we would need to analyze their meaning on a sentence-by-sentence basis, which was not feasible; thus, we left such ambiguous labels as-is.
 Glosses not in UniMorph
@@
 However, this would be an imprecise mapping, and could be confusing to a linguist of the particular language, so we again elect to leave these glosses unmapped.
-We use our mapping to normalize the dataset and will make the normalized version available in addition to the original.
+We use our mapping to normalize the dataset and make the normalized version available in addition to the original.
 4.3 Use in Future Research
@@
 NLP researchers benefit from a single, easily-accessible dataset covering many languages, which can be used for future research on interlinear glossing, morpheme analysis, segmentation, and translation.
-Linguistics researchers will be able to use the dataset to easily search for phenomena across languages, particular with the normalized version of the dataset.
+Linguistics researchers will be able to use the dataset to easily search for phenomena across languages, particularly with the normalized version of the dataset.
 For example, a linguist could find examples of sentences demonstrating the ergative/absolutive distinction.
-They could further refine this analysis by narrowing the results to related languages, using the glottocodes in our dataset.
+They could further refine this analysis by narrowing the results to a set of related languages, using the glottocodes in our dataset.
 As another example, if a linguist wished to determine how prior researchers have annotated examples in a particular language, they would previously have to search across research papers, textbooks, and small corpora.
@@
 Language Pre-train Train Eval Test
-Other languages 313,790 - - -
+Other languages 198,121 - - -
 In-domain languages
@@
 This could be valuable to real-world documentation projects, as we can serve a single pretrained model that can be used across projects, significantly reducing the barrier to using an automated system.
+The languages evaluated here are well-represented in the pretraining corpus, from Tsez (3.7k unsegmented examples, about 3% of the total corpus) to Arapaho (39.1k examples, 21%).
+A natural question is whether the model retains good performance on a language which occurs very rarely in the pretraining corpus.
+We simulate this scenario by adding a small amount of data to the pretraining corpus for two unseen languages: 1000 examples in Nyangbo and 500 in Natugu (less than 1% of the total corpus).
+We evaluate on the unseen test split and observe 76.8% and 55.0% morpheme accuracy, respectively.
+These results indicate the model can still perform well on languages that are sparse in the pretraining corpus.
 7 Results
@@
 The Tü-CL model (Girrbach, 2023a), which is a close second and outperforms on Gitksan and Lezgi, uses explicit latent segmentations, which seems to be particularly beneficial for the very low-resource, unsegmented setting.
+To illustrate common errors from the finetuned GlossLM moderls, we include examples of system outputs in Table 10.
+When inspecting outputs, we observe that there are sometimes inconsistencies in the IGT labels where multiple interchangeable glosses are used for the same morpheme.
+While we try to account for a portion of grammatical gloss variations (§4, §7.3), this is particularly an issue for lexical glosses (e.g.
+the Arapaho word ’eeneisih’i is glossed in the data as “how.X.things.are.named”, “how.s.o..is.named”, and “how.things.are.named”).
+We also find outputs that include lexical items present in the translation that are not included in the gold gloss, indicating that the model may rely too heavily on translations when predicting lexical glosses in certain cases.
 7.2 Comparison with Finetuned ByT5
@@
 A possible explanation is that even if there are no similar languages in the pretraining corpus, the GlossLMft can leverage knowledge about IGT formatting from unrelated languages.
-With enough data (starting with Nyangbo, 2100 examples), the two approaches achieve nearly identical performance.
+For Lezgi—which shows the greatest improvements from pretraining with the GlossLM corpus—a qualitative analysis of examples with the greatest morpheme error rate between finetuned ByT5 and GlossLM reveals that there are regular error patterns that are fixed with continual pretraining.
+For example, the finetuned ByT5 model often outputs AOC instead of AOR and OLB instead of ERG, whereas the finetuned GlossLM gets these correct.
+We include examples of these outputs in Table 11.
+However, with enough data (starting with Nyangbo, 2100 examples) the two approaches achieve nearly identical performance.
 This is an unsurprising result, indicating that large amounts of high-quality monolingual data overshadow any benefits from crosslingual transfer.
@@
 When comparing the performance of the pretrained model before finetuning on in-domain languages, we find minimal differences compared to the results in §6.3.
-This is also the case after finetuning, though we found improvements in three of the four out-of-domain languages: Lezgi ($+\Delta 2.9$), Natugu ($+\Delta 6.1$), and Nyangbo ($+\Delta 2.5$), which (aside from Gitksan) are the three lowest resourced languages in our test set, exceeding current SOTA performance.
-From this, it appears that normalization may be most helpful when adapting the pretrained model to an unseen language with a low-to-moderate amount of training data.
+Figure 6: Change in morpheme accuracy after normalizing glosses to the UniMorph schema and finetuning GlossLM. We observe small improvements for several languages, but worse performance in two cases.
+We report the change in morpheme accuracy after normalizing in Figure 6.
+We observe mixed results.
+Some languages (Arapaho, Uspanteko, Gitksan) show worse or equivalent performance.
+Others (Tsez, Nyangbo, Natugu, and Lezgi) show small to moderate improvements, with Lezgi achieving the largest improvement of 3.0 percentage points.
+Thus, we found that normalization was most helpful when finetuning the pretrained model on unseen languages with a low-to-moderate amount of training data.
 8 Related Works
@@
 The large IGT datasets for endangered languages (Arapaho, Guarani, Uspanteko) were collected in collaboration with native communities, and our work is in accordance with the agreements for their usage.
+Acknowledgements
+We would like to thank the three anonymous reviewers and meta-reviewer for their helpful feedback on this work.
+Parts of this work were supported by the National Science Foundation under Grant No.
+2149404, “CAREER: From One Language to Another” and Grant No.
+2211951, "From Acoustic Signal to Morphosyntactic Analysis in One End-to-End Neural System." Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
 Appendix A Language Distribution
-Figure 6: Counts per language.
-We only show languages with at least 2k samples present in the dataset.
-Arapaho (arap1274) is by far the most represented language in our data, followed by Uspanteko (uspa1245).
-Both languages are part of the SIGMORPHON Shared Task dataset.
-Figure 7: Counts per language family in our dataset.
-We only show language families with at least 1k samples present in the dataset.
-Language isolates and languages without recorded families in Glottolog Hammarström et al.
-(2023) are categorized under “None”.
-In Figure 6 and Figure 7, we visualize the number of examples per language and language family on a portion of our dataset.
+Figure 7 and Figure 8 display the number of examples per language and language family on a portion of our dataset.
 Appendix B Grambank Typological Analysis
@@
 Table 4: Pretraining Hyperparameters
-Appendix D Full Results
-We provide full results for accuracy and chrF++ scores in Table 5 and Table 6 . For our normalization experiments, we only trained and tested on unsegmented data for the target languages.
 Model Morpheme accuracy / Word accuracy (Segmented)
@@
 GlossLMall, pre refers to performance using the pretrained GlossLM directly, while GlossLMall, ft refers to performance after finetuning on the specific language.
+Appendix D Full Results
+We provide full results for accuracy and chrF++ scores in Table 5 and Table 6 . For our normalization experiments, we only trained and tested on unsegmented data for the target languages.
 Appendix E In- vs out-of-vocabulary errors
@@
 Top is the segmented setting (GlossLMall, ft), bottom is unsegmented (GlossLMunseg, ft).
-We report word-level accuracy for our finetuned GlossLM models, indicating whether the transcribed word is in- or out-of-vocabulary in Table 7, as well as the percent of OOV words in the test set.
-The OOV rate between segmented and unsegmented may vary slightly, as mappings between segmented and unsegmented forms are not necessarily one-to-one.
-We consider a word to be in-vocabulary if the form of the word in the transcription and its corresponding gold label in the gloss co-occur in the training data.
-We also include morpheme-OOV rates and statistics on lexical gloss overlap with translations in in Table 8 and Table 9, as reported in Ginn et al.
-(2023).
 Language Morpheme %OOV
@@
 Table 9: Percent of lexical glosses present in the translation in the test split. Nyangbo examples do not include translations.
-Appendix F Example System Outputs
+We report word-level accuracy for our finetuned GlossLM models, indicating whether the transcribed word is in- or out-of-vocabulary in Table 7, as well as the percent of OOV words in the test set.
+The OOV rate between segmented and unsegmented may vary slightly, as mappings between segmented and unsegmented forms are not necessarily one-to-one.
+We consider a word to be in-vocabulary if the form of the word in the transcription and its corresponding gold label in the gloss co-occur in the training data.
+We also include morpheme-OOV rates and statistics on lexical gloss overlap with translations in in Table 8 and Table 9, as reported in Ginn et al.
+(2023).
+Appendix F Example Outputs
+We include example outputs to show the errors discussed in §7.1 and §7.2.
 Sample ID
@@
 output himself- FOC be- PERF- PTP real.story- PL was
-lezg1247_1 transcription ва гьада гьикъван гьа терези- ди- н стрелка пара хкаж хьа- нва- тІа гьам вилик кутуна- нва
+lezg1247_22 transcription ва гьада гьикъван гьа терези- ди- н стрелка пара хкаж хьа- нва- тІа гьам вилик кутуна- нва
 translation she put first that person which is more valuable according to the position of scale arrow.
@@
 Table 10: Selected example outputs to illustrate errors by GlossLMall finetuned models.
-We include examples of system outputs to illustrate the types of errors discussion in Table 10.
+Sample ID MER
+lezg1247_71 transcription Акъадарда и пачагь ибуру ламрал , балкІандал акъадарда , ламрал акъадарда яда , цІайни ахъайда , им гьи хуьрей агъуз .
+gold mount.ENT this king these-ERG donkey.on horse-SPSS mount.ENT donkey.on mount.ENT or fire.and released.ENT he this village.out.of down
+ByT5 mount.ENT-ENT this king these-ERG donkey-OBL-SPSS horse-OBL-SPSS mount.ENT-ENT donkey-OBL-SPSS mount.ENT-ENT was fire-FOC 0.585
+   put.ENT-ENT he this village-INESS-SPSS down.ENT-ENT
+GlossLM mount.ENT this king these-ERG donkey.on horse-SPSS mount.ENT donkey.on mount.ENT or fire.and mount.ENT he this village.out.of down 0.068
+lezg1247_49 transcription Хтана балкІан , « Гьаа », лагьана « гила чавай физ жеда » лагьана , « пачагьдин руш гъиз ».
+gold return-AOR horse Yes say-AOR now help-INELAT go-DAT be-ENT say-AOR king-DIR-GEN girl will.bring-DAT
+ByT5 coming-AOC horse Yes say-AOC now king-DIR-GEN girl be-ENT say-AOC king-DIR-GEN girl be-ENT 0.398
+GlossLM return-AOR horse Yes say-AOR now help-INELAT go-DAT be-ENT say-AOR king-DIR-GEN girl will.bring-DAT 0
+lezg1247_27 transcription АтІуз гана ибурун кьилерни .
+gold cutting.IMC-DAT give-AOR these-ERG-GEN chapter-PL-FOC
+ByT5 then give-AOR these-ERG-GEN head-PL-FOC 0.380
+GlossLM cutting.IMC-DAT give-AOR these-ERG-GEN chapter-PL-FOC 0
+lezg1247_0 transcription И гададини гъил вегьена са жуьт къачуда .
+gold this boy-DIR-GEN-ERG hand threw-AOR one pair take-INESS
+ByT5 this boy-DIR-GEN-ERG hand took-AOC one necklace take-ENT 0.327
+GlossLM this boy-DIR-GEN-ERG hand threw-AOR one pair take-INESS 0
+lezg1247_42 transcription И рушни пачагь хьана гила башламишда вичин пачагьвализ .
+gold this girl-Q king happened-AOR now started.ENT-INESS himself-ERG-GEN reigned.ENT-ERG-DAT
+ByT5 this girl-FOC king be-AOC now started.ENT-ENT himself-OBL-GEN reign.to-OBL-DAT 0.325
+GlossLM this girl-Q king happened-AOR now started.ENT-INESS himself-ERG-GEN reigned.ENT-ERG-DAT 0
+Table 11: Lezgi examples with the highest difference in MER between finetuned ByT5 and GlossLM outputs.
 ID Avg. Value ID (cont.) Avg. Value (cont.) ID (cont.) Avg. Value (cont.)